### v3

# Predicting the Distribution of Treatment

Abstract
Important questions for impact evaluation require knowledge not only of average effects, but of the distribution of treatment effects. The inability to observe individual counterfactuals makes answering these empirical questions challenging. I propose an inference approach for points of the distribution of treatment effects by incorporating predicted counterfactuals through covariate adjustment. I provide finite-sample valid inference using sample-splitting, and asymptotically valid inference using cross-fitting, under arguably weak conditions. Revisiting five randomized controlled trials on microcredit that reported null average effects, I find important distributional impacts, with some individuals helped and others harmed by the increased credit access.
Keywords: Distribution of treatment effects, heterogeneous treatment effects, machine learning, causal inference, microcredit.
JEL classification codes: C12, C14, O12.
1 Introduction
Important questions in impact evaluation cannot be answered by average treatment effects alone. What proportion of individuals are harmed by a treatment? Does a policy help many by a little? Or a few by a lot? The distribution of treatment effects carries crucial information with important equity and efficiency implications. For example, if a policy with positive average effect is found to harm a significant portion of the population, it may be crucial to prioritize mitigating the harm or reconsider its implementation. If a few individuals have large benefits, it may be important to understand why and how to efficiently target the policy recipients. However, learning the distribution of treatment effects is challenging because of the fundamental problem of causal inference: one can never observe individual counterfactuals. But one may be able to predict them.
The main contribution of this paper is to provide a new approach to inference on points of the distribution of treatment effects,
$\theta=P(Y(1)-Y(0)\leq\delta),$
for any value of $\delta$, by leveraging covariate-adjustment and predictive models. The setting assumes a binary treatment $D$ and that the researcher has access to a representative sample of $Y(1)$ and $Y(0)$, the potential outcomes under treatment and no treatment, respectively, and pre-treatment covariates $X$ [^1] In Online Appendix OA.1, I provide an alternative setting with a known propensity score. This is typically the setting in Randomized Controlled Trials (RCTs). I denote $\theta$ by Distributional Treatment Effect (DTE) and the method by CAIDE (Covariate Adjustment for Inference on Distributional Effects). All results are pointwise in $\delta$.
Relative to previous literature, my first contribution is to provide a new characterization of the identified set of $\theta$ which is more amenable to inference. Based on this result, I develop two novel methods for inference. The first is valid in finite samples and requires arguably weak conditions. For example, the conditions are satisfied if the data come from an RCT. To my knowledge, this is the first finite-sample inference result on the DTE that incorporates covariates. The second inference method is asymptotically valid and more powerful than the finite-sample approach. Both inference methods rely on estimating a covariate-adjustment term, and I show they are robust to its misspecification, in the sense that a misspecified term leads to a valid identified set, although wider. If the covariate-adjustment term is correctly specified, the identified set is sharp. As a consequence of this robustness, the covariate-adjustment term can be estimated using any method, including machine learning algorithms. The relevance of my contribution is demonstrated in an application to microcredit, where I find important distributional impacts, specifically statistically significant bounds on the proportion of individuals helped and hurt by increased credit access. While academics and policymakers have hypothesized and debated the theory underlying such impacts, their empirical estimation has been hampered by methodological limitations.
My results can be contrasted with those in Fan and Park (2010), who characterized the sharp identified set of $\theta$ in the presence of covariates. Their approach to inference focuses on the case without covariates since it is challenging to implement inference based on their characterization of the identified set, as discussed in Remark 1. In contrast, my characterization of the sharp identified set is more amenable to inference and satisfies the robustness property mentioned above.
Asymptotic inference on $\theta$ in the presence of covariates was studied by two recent and independent working papers, Semenova (2023) and Ji et al. (2023). Both papers consider the more general problem of learning averages of intersection bounds, which includes the DTE as a particular case. Their approach is based on the dual of an optimization problem, and the confidence intervals of both papers coincide for the DTE. By leveraging the specific structure of the problem of learning the DTE, my main advantage relative to these alternatives is that CAIDE always estimates bounds that are (weakly) narrower, leading to more power under suitable conditions (Appendix A.1). In practice, this difference can lead to different conclusions regarding statistical significance, as shown in a Monte Carlo study in Section 5 and in the application in Section 6. Another difference relative to the two papers is that I allow for using the proportion of treated units in-sample instead of the true propensity score, which further reduces the confidence interval length (Appendix A.2). Finally, I show asymptotic exactness under continuous outcomes, while Ji et al. (2023) restrict the support of the outcomes to a finite set to derive exactness. Semenova (2023) and Ji et al. (2023) do not provide finite-sample inference.
To understand my approach, consider a scenario where a researcher aims to predict $Y(0)$ using a function of covariates $s(X)$. An example of such a function is the conditional expectation, $s(x)=\mathbb{E}[Y(0)|X=x]$. Ideally, $s(X)$ would be close to $Y(0)$, and the distribution of $Y(1)-s(X)$ would approximate the distribution of treatment effects. However, $s(X)$ is generally an imperfect predictor for two main reasons. First, because $Y(0)$ is rarely a deterministic function of observable covariates, so one can typically only approximately predict it. Second, because $s(X)$ needs to be calculated from the data, and is subject to random sampling error. CAIDE is designed to address both of these challenges. In this example, the idea is that even if $s(X)$ is an imperfect predictor, examining the distribution of $Y(0)-s(X)$ (the prediction error) can help correct for the difference between $Y(1)-s(X)$ and the actual treatment effect. Intuitively, if $|Y(0)-s(X)|$ is small (e.g., less than 0.1), the error between $Y(1)-s(X)$ and the treatment effect is also small (at most $0.1$). In general, correcting for prediction error using the entire distribution of $Y(0)-s(X)$ provides more information than simply using an upper bound.
The intuition above does not rely on the form of $s$. In fact, by combining the information from the distributions of both $Y(1)-s(X)$ and $Y(0)-s(X)$, any function $s$ can be used to learn $\theta$. However, when correcting for prediction error, different choices of $s$ can lead to different approximations of the distribution of treatment effects. CAIDE proceeds by estimating an optimal form of $s$, defined in Theorem 1, which depends on the distributions of both potential outcomes conditional on covariates — that is, it is not the prediction of just one potential outcome. In this case, the function $s$ can be interpreted as a covariate-adjustment term, since the same function is subtracted from both $Y(1)$ and $Y(0)$. In Section 2, I show that the optimal function is different for learning the lower or upper bound on $\theta$, and the procedure can be repeated with different functions $s_{L}^{*}$ and $s_{U}^{*}$ when one is interested in two-sided inference.
The fact that any function $s$ is valid motivates a sample-splitting strategy: one portion of the sample is used to estimate $s^{*}$ ($s^{*}_{L}$ or $s^{*}_{U}$) with some estimator $\hat{s}$ using any machine learning algorithm; in the remaining sample, I calculate a confidence interval for $\theta$ treating $\hat{s}$ as fixed. Once this function is fixed, estimation becomes a low-dimensional problem, similar to Fan and Park (2010), and valid inference can be achieved regardless of the statistical properties of $\hat{s}$. In fact, finite-sample inference on $\theta$ (Theorem 2) is valid when data come from an RCT, without requiring further regularity conditions. Moreover, cross-fitting, a more efficient way of sample-splitting, can be used for asymptotic inference under suitable conditions (Theorem 4). In both cases, if $\hat{s}$ is a poor estimate of $s^{*}$, the resulting confidence interval is wider but still valid compared to an estimator $\hat{s}$ that is close to $s^{*}$. I compare both approaches to inference in Section 5 through a Monte Carlo study.
I apply CAIDE to study heterogeneous effects of microcredit (Section 6). Microcredit has reached more than 170 million borrowers worldwide in 2022 (Convergences, 2023) and has been the subject of hundreds of academic papers, including several RCTs. Theory suggests that microborrowing can be either beneficial or detrimental (Banerjee, 2013, Garz et al., 2021), and most empirical evidence indicates that average effects are in general mild if not negligible (Banerjee et al., 2015b, Meager, 2019). Analyses of heterogeneous effects have mostly been limited to Quantile Treatment Effects (QTEs), suggesting mainly positive impacts for some and rarely negative impacts (Meager, 2022). I revisit five studies that randomized microloans at the individual level and find evidence of heterogeneity, including negative treatment effects for some households. The estimated lower bound on the proportion of individuals who would be better off without the loan varies from $2.8\%$ to $26.0\%$, and the lower bound on the proportion who benefited ranges from $2.9\%$ to $26.1\%$.
The paper proceeds as follows. Section 1.1 reviews the related literature. Section 2 describes the setup and shows the main identification results of the paper. Section 3 provides finite-sample inference, and Section 4 asymptotic inference and exactness using cross-fitting. Section 5 presents a Monte Carlo study, and Section 6 the application to microcredit. Finally, Section 7 concludes. Proofs not displayed in the main text are deferred to the Appendix.
1.1 Related Literature
Early work on bounding the distribution of treatment effects are Heckman et al. (1997) and Manski (1997), who propose assumptions that restrict the joint distribution of potential outcomes in order to achieve informative bounds on $\theta$. Makarov (1982) and Williamson and Downs (1990) characterize the sharp bounds on $\theta$ given only knowledge of the marginal distributions of $Y(1)$ and $Y(0)$. [^2] Both papers consider the generic problem of bounding the distribution of the sum of two random variables given only knowledge of their marginals. Fan and Park (2010) use these bounds to provide asymptotic inference on $\theta$, and Ruiz and Padilla (2022) provide finite-sample inference.
Fan and Park (2010) also characterize sharp bounds on $\theta$ in the presence of pre-treatment covariates, but they do not approach inference based on these bounds. It is challenging to provide inference from their characterization, as discussed in Remark 1. In contrast, by incorporating covariates via a covariate-adjustment term, the characterization of the sharp bounds that I provide is amenable to inference.
Kallus (2022) studies the sharp bounds in the presence of covariates in the case of binary outcomes and proposes inference under high-level assumptions on the data-generating process. Beyond the binary case, two recent and independent working papers studied inference using the sharp bounds on $\theta$ in the presence of covariates, Semenova (2023) and Ji et al. (2023). The two papers study the more general problem of averages of intersection bounds, which includes $\theta$ as a particular case. Semenova (2023) provides inference when the conditional distributions of potential outcomes can be estimated uniformly at the $n^{-1/4}$ rate and potential outcomes are discrete. Ji et al. (2023) uses optimal transport theory to show that solving the dual leads to valid bounds even if conditional distributions are misspecified. In contrast, my approach is based on covariate adjustment. The main differences are that I provide finite-sample inference, asymptotic inference that is more powerful under suitable conditions (Appendix A.1), a further reduction in the length of confidence intervals by using the proportion of treated units instead of the propensity score (Appendix A.2), and asymptotic exactness under continuous outcomes.
This paper also contributes to the literature on using machine learning for estimating heterogeneous treatment effects (Athey and Imbens, 2016; Chernozhukov et al., 2018; Wager and Athey, 2018; Athey et al., 2019; Yadlowsky et al., 2021; Knaus et al., 2022; Athey et al., 2023; Chernozhukov et al., 2023). Most papers in this literature focus on estimating the point-identified conditional average treatment effect (CATE) and use the estimated CATE to predict individual treatment effects. CAIDE is complementary to these approaches. I focus on the partially-identified DTE $\theta$, which answers a different type of question. For example, if the CATE is zero for all values of $X$, many individuals could still be harmed, and many could benefit. In this case, the DTE could reveal, for example, the proportion of individuals who are harmed and who benefit.
This paper is also related to the literature on conformal prediction (Vovk et al., 2005; Lei et al., 2018). Lei and Candès (2021) propose a prediction interval to the individual treatment effect, i.e., an interval to cover the random variable $Y(1)-Y(0)$ for a specific individual with a certain probability. In contrast, I focus on the population parameter $\theta=P(Y(1)-Y(0)\leq\delta)$. Also related is the literature on covariate adjustment in the context of estimating the Average Treatment Effect (ATE) in RCTs (e.g., Lin, 2013, Wager et al., 2016, Rothe, 2018, Wu and Gagnon-Bartsch, 2018; review in Van Lancker et al., 2023; comparisons in Kahan et al., 2016, Tackney et al., 2023). Most of this literature focuses on adjusting outcomes for covariates to improve precision. I use covariate adjustment to improve identification (i.e., to get narrower bounds) of a different estimand, the DTE.
I add to the development economics literature on evaluating the impacts of microcredit. Metanalyses suggest that average impacts are small if not negligible [^3] Breza and Kinnan (2021) find negative effects of the shutdown of microfinance institutions in Andhra Pradesh, which suggests that microcredit might have large and positive equilibrium effects. (Banerjee et al., 2015b, Meager, 2019), however, there is growing evidence that impact is heterogeneous (Banerjee et al., 2019, Bryan et al., 2021, Meager, 2022, Cueva et al., 2024). Analysis of heterogeneity suggested mostly positive impact in the upper tail of income/profits distributions (Angelucci et al., 2015, Augsburg et al., 2015, Banerjee et al., 2015a, Meager, 2022). Yet, evidence of negative effects has been scarce [^4] Crépon et al. (2015), for example, finds negative quantile treatment effect on profits at the 10th percentile. However, they interpret the evidence as a “reduced form”: “they do not necessarily mean that the impact of getting credit itself has the same heterogeneity”, since take-up is low (randomization was conducted at the village level) and “we do not know where the compliers lie in the distribution of outcomes”. They also suggest the negative effect “might be partially due to long-term investments misclassified as current expenses”, supported by a positive estimated QTE on consumption at the 10th percentile.. For example, through a metanalysis of seven studies, Meager (2022) finds “no generalizable negative quantile treatment effects”. I revisit five RCTs in microcredit and find novel evidence of both positive and negative effects.
2 Setup and Identification via Covariate Adjustment
In this section, I define the setup of the paper and present the identification results of $\theta=P(Y(1)-Y(0)\leq\delta)$ that motivate the confidence intervals proposed in Sections 3 and 4. There are two sources of ambiguity when learning $\theta$ from a sample and predicted counterfactuals. The first is the usual statistical uncertainty due to random sampling error, the usual uncertainty in estimating a population parameter from a sample. The second is due to the prediction of the counterfactuals. Even with a large sample size, the predicted counterfactuals can generally approximate the true counterfactuals, but they do so with error. The consequence of this ambiguity is that, even with access to population data, in general, one can at best find an interval (or bounds) that contain $\theta$ (see, e.g., Heckman et al., 1997; Manski, 1997).
This section lays out the strategy I use to deal with the latter. The first result is a new identifying equation for $\theta$ in (2.3), that is, an interval, denoted identified set, that contains $\theta$ expressed in terms of the distribution of the observable data. This equation incorporates the information on predicted counterfactuals through a scalar function $s(x)$, which acts as a covariate-adjustment term. The main benefit of (2.3) is that the interval contains $\theta$ for any function $s$. That is, (2.3) might be an outer identified set, but it is always a valid identified set. This property is crucial so that the confidence intervals proposed in Sections 3 and 4 are valid regardless of the statistical properties of how $s$ is estimated.
The second result is Theorem 1, which characterizes optimal forms of $s$ that make the identified set in (2.3) sharp, that is, the narrowest possible interval that has the guarantee to contain $\theta$. As it turns out, in general, no single choice of $s$ makes both lower and upper bounds sharp; instead, each bound requires a different function $s$, denoted $s^{*}_{L}$ and $s^{*}_{U}$ respectively for the lower and upper bounds. The sharp identified set in Theorem 1 is an alternative characterization to the one provided in Fan and Park (2010). I argue in Remark 1 that my representation leads to two major advantages for inference. Most importantly, I argue that inference using my characterization is tractable under arguably mild conditions (Theorems 2 and 4) when the bounds are estimated nonparametrically. In contrast, I show that it is hard to characterize the asymptotic distribution of the plug-in estimator based on Fan and Park (2010)’s characterization, even when using sample-splitting, when the bounds are estimated nonparametrically. Note that Fan and Park (2010) do not address inference on $\theta$ in the presence of covariates.
Finally, in subsection 2.1 I discuss cases when $\theta$ is point-identified, which happens when covariates are informative enough to predict the sign of the difference of counterfactuals.
Without loss of generality and to simplify notation, I focus on $\delta=0$. For a different value of $\delta$, all results hold by redefining $\tilde{Y}(0)=Y(0)+\delta$. Consider the standard potential outcomes setup, where $Y\in\mathcal{Y}\subseteq\mathbb{R}$ is the outcome, $D\in\{0,1\}$ is the treatment assignment indicator, $Y(1)$ and $Y(0)$ are potential outcomes connected by $Y=DY(1)+(1-D)Y(0)$, and $X\in\mathcal{X}$ denotes a set of pre-treatment covariates. The setting is formalized in Assumption 1.
Assumption 1.
Let $(Y,D,X)\sim P$. For some $Y_{\ell},Y_{u}\in\mathbb{R}$ and $B>0$, $P$ satisfies:
  1. (i)
    $Y\in[Y_{\ell},Y_{u}]$;
  2. (ii)
    $(Y(1),Y(0))\perp D|X$; the propensity score $p(x)=P(D=1|X=x)=\pi$ does not depend on $x$ and satisfies $B<\pi<1-B$.
Assumption A1(i) of bounded $\mathcal{Y}$ is technical, and is discussed in Remark 2. A1(ii) is the usual unconfoundedness assumption, and it is used for identification of the marginals of $Y(1)$ and $Y(0)$. It typically holds, for example, in Randomized Controlled Trials (RCTs), where $D$ is randomized. The constant propensity score assumption facilitates exposition. In the asymptotic analysis, this assumption can be relaxed to allow for an arbitrary, known propensity score, as demonstrated in Online Appendix OA.1. Extending my results to accommodate unknown propensity scores that must be estimated from the data is an interesting avenue for future research.
The starting point for CAIDE is the fact that for any measurable function $s:\mathcal{X}\rightarrow\mathcal{Y}$, it holds that
$\theta=Pr\{Y(1)-Y(0)\leq 0\}=Pr\{[Y(1)-s(X)]-[Y(0)-s(X)]\leq 0\}.$ (2.1)
Therefore, I propose to bound $\theta$ using the marginal distributions of covariate-adjusted outcomes $Y(1)-s(X)$ and $Y(0)-s(X)$. This strategy is in contrast to Fan and Park (2010), which considered the marginals of $Y(1)$ and $Y(0)$ for inference. A major advantage of my approach is that it is flexible in the covariate adjustment term, in the sense that any $s$ induces bounds that contain $\theta$. Such flexibility is crucial to ensure that the estimation of $s$, regardless of the method or number of covariates used, leads to valid inference in Sections 3 and 4. The Makarov bounds (Makarov, 1982; Williamson and Downs, 1990), used in Fan and Park (2010), always contain $\theta$:
$\theta\in\left[\max_{t}\{F_{1}(t)-F_{0}(t)\},1+\min_{t}\{F_{1}(t)-F_{0}(t)\}\right],$ (2.2)
where $F_{j}(t)=P(Y(j)\leq t)$, $j\in\{0,1\}$. Hence, by replacing the random variables $Y(1)$ and $Y(0)$ by $Y(1)-s(X)$ and $Y(0)-s(X)$, the bounds induced by $s$ contain $\theta$:
[…]
where
$\Delta(t,s)=P(Y(1)-s(X)\leq t)-P(Y(0)-s(X)\leq t).$ (2.4)
Because of the equivalence in (2.1), and the validity of the Makarov bounds (2.2), (2.3) contains $\theta$ for any $s$. If covariates are not used, e.g. if $s(x)=0$, then (2.3) is equivalent to (2.2). Different functions $s$ lead to different bounds, with varying lengths. For specific functions defined in Theorem 1, the induced bounds are sharp, that is, the smallest possible valid bounds given the information from the data. Fan and Park (2010) derived one characterization for these bounds:
$[\theta_{L}^{*},\theta_{U}^{*}]=\left[\mathbb{E}\left[\max_{t}\{F_{1}(t|X)-F_{0}(t|X)\}\right],1+\mathbb{E}\left[\min_{t}\{F_{1}(t|X)-F_{0}(t|X)\}\right]\right],$ (2.5)
where $F_{j}(t|x)=P(Y(j)\leq t|X=x)$, $j=0,1$. In Remark 1, I compare (2.5) and Theorem 1.
Theorem 1 (Sharp transformations $s^{*}_{L}$ and $s^{*}_{U}$).
Let Assumption 1(ii) hold. Define
$s^{*}_{L}(x)$ $\in\arg\max_{t}\left\{F_{1}(t|x)-F_{0}(t|x)\right\},$ (2.6)
$s^{*}_{U}(x)$ $\in\arg\min_{t}\left\{F_{1}(t|x)-F_{0}(t|x)\right\}.$ (2.7)
Then, $\left[\max_{t}\Delta(t,s^{*}_{L}),1+\min_{t}\Delta(t,s^{*}_{U})\right]=[\theta_{L}^{*},\theta_{U}^{*}]$ are the smallest valid (sharp) bounds on $\theta$ given the distribution of the observable data $(Y,D,X)$. That is, for every $\theta$ in this interval, there exists a joint distribution of $(Y(1),Y(0),D,X)$ with the same distribution of the observable data such that $P(Y(1)-Y(0)\leq 0)=\theta$.
Proof.
For the lower bound,
$\theta_{L}^{*}$ $\geq\max_{t}\Delta(t,s^{*}_{L})$ (2.8)
$\geq\Delta(0,s^{*}_{L})$
$=Pr\{Y(1)-s^{*}_{L}(X)\leq 0\}-Pr\{Y(0)-s^{*}_{L}(X)\leq 0\}$ (2.9)
$=\mathbb{E}\left[Pr\{Y(1)-s^{*}_{L}(X)\leq 0|X\}-Pr\{Y(0)-s^{*}_{L}(X)\leq 0|X\}\right]$ (2.10)
$=\mathbb{E}\left[F_{1}(s^{*}_{L}(X)|X)-F_{0}(s^{*}_{L}(X)|X)\right]$ (2.11)
$=\mathbb{E}\left[\max_{t}\left\{F_{1}(t|X)-F_{0}(t|X)\right\}\right]$ (2.12)
$=\theta_{L}^{*}$ (2.13)
where (2.8) holds from (2.3) and since $\theta_{L}^{*}$ is sharp (Fan and Park, 2010); (2.9) by definition of $\Delta$; (2.10) by the law of iterated expectations; (2.11) by definition of $F_{j}(t|x)$ ($j=0,1$); (2.12) by the definition of $s^{*}_{L}$; and (2.13) by definition of $\theta_{L}^{*}$ (2.5). The proof for the upper bound is analogous. ∎
The result of Theorem 1 is new. It shows that the sharp bounds on $\theta$ in the presence of covariates can be obtained via covariate adjustment. That is, for specific functions $s^{*}_{L}(x)$ and $s^{*}_{U}(x)$, the induced bounds in (2.3) are equal to the sharp bounds in (2.5). Note that $s^{*}_{L}(x)$ and $s^{*}_{U}(x)$ need not be unique and are generally different. Hence, in general, there is no single function $s$ that induces both lower and upper bounds to be sharp in (2.3), and in Sections 3 and 4 I propose to estimate both $s^{*}_{L}(x)$ and $s^{*}_{U}(x)$.
Remark 1.
I argue that my representation of the identified set in Theorem 1 has two major advantages for inference compared to Fan and Park (2010)’s characterization in (2.5).
The first advantage is that Theorem 1 enables a tractable approach to inference under arguably mild conditions when the conditional distributions are estimated nonparametrically (Theorems 2 and 4). In contrast, deriving the asymptotic distribution of the plug-in estimator based on Fan and Park (2010)’s characterization is challenging, even when using sample-splitting, when the conditional distributions are estimated nonparametrically. I show in Appendix A.3 that this asymptotic distribution depends on the specific estimation method used for the conditional distributions, and the convergence rate of the plug-in estimator depends on how quickly they are learned. It is difficult to characterize this asymptotic distribution without an assumption on the convergence rates of the conditional distribution estimators. However, such assumption often does not hold when the dataset includes more than a few covariates and the conditional distributions are estimated nonparametrically.
The second advantage of my identifying equation is that the bounds in (2.3) are robust to misspecification of the covariate-adjustment terms when they are learned parametrically. That is, if $(s^{*}_{L},s^{*}_{U})$ are estimated inconsistently, the limit bound may be wider than the sharp bounds, but they always contain $\theta$. As a consequence, the inference results established in Theorems 2 and 4 are valid regardless of whether $(s^{*}_{L},s^{*}_{U})$ are estimated nonparametrically or with a potentially misspecified parametric model. In contrast, the plug-in estimator based on Fan and Park (2010)’s characterization can converge to bounds narrower than the sharp, and thus not contain $\theta$, if the conditional distributions are estimated with a misspecified parametric model.
Moreover, this robustness property is also advantageous when $(s^{*}_{L},s^{*}_{U})$ are estimated consistently, since noisy estimates of the covariate-adjustment terms translate into estimating bounds that are wider. In contrast, the plug-in estimator based on Fan and Park (2010)’s characterization does not have this property. This robustness property is relevant both for the finite-sample valid inference approach of Theorem 2, which relies only on an RCT assumption, and the large-sample approach of Theorem 4.
Remark 2.
The assumption of bounded $\mathcal{Y}$ ensures that $s^{*}_{L}(x)$ and $s^{*}_{U}(x)$ are finite, which is important for the equivalence in (2.1) when using the marginals of $Y(1)-s(X)$ and $Y(0)-s(X)$. This assumption could easily be relaxed by defining a new outcome $\widetilde{Y}=\Phi(Y)$, where $\Phi$ is any bounded, strictly increasing function, such as the standard normal cdf. $\widetilde{Y}$ is always bounded, and $Pr\{Y(1)-Y(0)\leq 0\}=Pr\{\widetilde{Y}(1)-\widetilde{Y}(0)\leq 0\}=Pr\{[\widetilde{Y}(1)-s(X)]-[\widetilde{Y}(0)-s(X)]\leq 0\}$ holds.
Remark 3.
In principle, functions $s$ that differ from the optimal $s^{*}_{L}(x)$ and $s^{*}_{U}(x)$ can lead to wider bounds, potentially even wider than those obtained using $s(x)=0$. This phenomenon, however, is not of concern when $s^{*}_{L}(x)$ and $s^{*}_{U}(x)$ are estimated nonparametrically, since they yield the sharp bounds, which are by construction (weakly) narrower than those using $s(x)=0$.
2.1 Conditions for Point Identification
Blue lines represent the distribution of $Y(1)$ conditional on a given value of $X=x$, and red lines the conditional distribution of $Y(0)$. Panels in the first row show the cumulative distributions, and in the second row, the probability density functions. The left column shows a scenario with partial identification, and the right column shows a scenario with point identification. The green dashed line represents the point $t$ that maximizes $F_{1}(t|x)-F_{0}(t|x)$. In both scenarios $\min_{t}F_{1}(t|x)-F_{0}(t|x)=0$, for example for $t=0$.
Figure 1: Example scenarios of distributions of potential outcomes conditional on a given value of covariates $X=x$.
From equation (2.5), a sufficient condition for point identification of $\theta$ is that $\max_{t}\{F_{1}(t|x)-F_{0}(t|x)\}=1+\min_{t}\{F_{1}(t|x)-F_{0}(t|x)\}$ for all $x$. Figure 1 gives intuition for when this condition holds, focusing on a single value of $X=x$ for ease of exposition. All panels show the conditional distributions of $Y(1)$ (in blue) and $Y(0)$ (in red), which represent the prediction of the potential outcomes given $X=x$. The upper panels show the cumulative distribution and the lower panels show the probability density. Note that predictions are not a single point since potential outcomes are random variables even conditional on $X=x$. In both scenarios $\min_{t}\{F_{1}(t|x)-F_{0}(t|x)\}=0$, for example for $t=0$, and $F_{1}(t|x)-F_{0}(t|x)$ is maximized at $t=0.75$ (green dashed line).
In the lower left panel, the distributions of $Y(1)$ and $Y(0)$ are partially overlapping, and $\max_{t}\{F_{1}(t|x)-F_{0}(t|x)\}<1+\min_{t}\{F_{1}(t|x)-F_{0}(t|x)\}=1$ (upper left panel). Since there is overlap, one cannot accurately predict if $Y(1)-Y(0)<0$ or not. For example, if an individual is at the right tail of $Y(1)$ (blue) but at the left tail of $Y(0)$ (red), then $Y(1)-Y(0)>0$. If the opposite happens, then $Y(1)-Y(0)<0$. However, since the overlap is small, the probability that $Y(1)-Y(0)<0$ is large, and the fraction harmed (for this value of $X=x$) is close to point identified. In the lower right panel, the distributions are disjoint, and $\max_{t}\{F_{1}(t|x)-F_{0}(t|x)\}=1+\min_{t}\{F_{1}(t|x)-F_{0}(t|x)\}=1$ (upper right panel). In this case, even though the covariates are not informative enough to get a point prediction of $Y(1)$ and $Y(0)$, it is informative enough to learn that $Y(1)<Y(0)$ for this value of $X=x$, and thus all individuals with $X=x$ are harmed.
Note that the interval in (2.5) is an average of $\max_{t}\{F_{1}(t|x)-F_{0}(t|x)\}$ and $1+\min_{t}\{F_{1}(t|x)-F_{0}(t|x)\}$ over all values of $X$. Hence, the figure illustrates that the informativeness of covariates in predicting the outcomes is crucial for having a narrow interval for $\theta$. In fact, if the predictions (conditional distributions) of $Y(1)$ and $Y(0)$ are disjoint for all values of $X$, then $\theta$ is point-identified. Note that in Figure 1, the green line represents the value of $s^{*}_{L}(x)$, which is the point $t$ that maximizes the distance $F_{1}(t|x)-F_{0}(t|x)$. Hence, $s^{*}_{L}(x)$ is the point that makes the two distributions as separated as possible. This fact is what motivates the notation $s$, for separator function.
3 Finite Sample Inference with Sample-Splitting
In this section, I propose sample-splitting estimators for $(\theta_{L}^{*},\theta_{U}^{*})$ and a confidence interval (CI) for $\theta$ that provides a finite sample guarantee of coverage. Given an iid sample $(Y_{i},D_{i},X_{i})_{i=1}^{n}\sim P$ that satisfies A1, the estimation strategy consists of randomly splitting the sample into two, using the first set to estimate $(s^{*}_{L},s^{*}_{U})$ with $(\hat{s}_{L},\hat{s}_{U})$, and the second set to estimate the bounds $[\max_{t}\Delta(t,\hat{s}_{L}),1+\min_{t}\Delta(t,\hat{s}_{U})]$ (as in 2.3) with a sample analogue. Conditional on the first sample, these bounds are valid due to (2.1), and finite sample inference is achieved by applying the DKW inequality (Dvoretzky et al., 1956; Massart, 1990).
I propose sample-splitting estimators for the sharp bounds $(\theta_{L}^{*},\theta_{U}^{*})$:
Definition 1 (Sample-splitting estimators).
  • (i)
    Randomly split the sample into two sets, denoted main and auxiliary sets. Denote by $\mathcal{I}_{M}^{1}\subset\{1,\dots,n\}$ the units in the main sample assigned to treatment, and $\mathcal{I}_{M}^{0}$ the units assigned to control;
  • (ii)
    Using data from the auxiliary sample only, estimate $(s^{*}_{L},s^{*}_{U})$ by replacing $F_{1}(t|X)$ and $F_{0}(t|X)$ in (2.6) and (2.7) by any estimators of the conditional cdfs $\widehat{F}_{1}(t|X)$ and $\widehat{F}_{0}(t|X)$. Denote the estimators $(\hat{s}_{L},\hat{s}_{U})$.
  • (iii)
    In the main sample, compute the final estimators
    $\widetilde{\theta}_{L}$ $=\max_{t}\left\{\frac{1}{|\mathcal{I}^{1}_{M}|}\sum_{i\in\mathcal{I}^{1}_{M}}\mathbb{I}\left(Y_{i}-\hat{s}_{L}(X_{i})\leq t\right)-\frac{1}{|\mathcal{I}^{0}_{M}|}\sum_{i\in\mathcal{I}^{0}_{M}}\mathbb{I}\left(Y_{i}-\hat{s}_{L}(X_{i})\leq t\right)\right\},$
    $\widetilde{\theta}_{U}$ $=1+\min_{t}\left\{\frac{1}{|\mathcal{I}^{1}_{M}|}\sum_{i\in\mathcal{I}^{1}_{M}}\mathbb{I}\left(Y_{i}-\hat{s}_{U}(X_{i})\leq t\right)-\frac{1}{|\mathcal{I}^{0}_{M}|}\sum_{i\in\mathcal{I}^{0}_{M}}\mathbb{I}\left(Y_{i}-\hat{s}_{U}(X_{i})\leq t\right)\right\}.$
$F_{1}(t|X)$ and $F_{0}(t|X)$ can be estimated with a variety of methods, including machine learning algorithms or other nonparametric estimators (see, e.g., Kneib et al., 2023 for a review). I discuss guidelines for estimating $\hat{s}$ in Appendix A.4. Although any method can be used to estimate $F_{1}(t|X)$ and $F_{0}(t|X)$, estimates $(\hat{s}_{L},\hat{s}_{U})$ closer to $(s^{*}_{L},s^{*}_{U})$ yield bounds $[\max_{t}\Delta(t,\hat{s}_{L}),1+\min_{t}\Delta(t,\hat{s}_{U})]$ that are closer to the sharp $[\max_{t}\Delta(t,s^{*}_{L}),1+\min_{t}\Delta(t,s^{*}_{U})]$, and poor estimators can lead to wider intervals.
A default choice is to make the main and auxiliary samples roughly the same size, but that is unnecessary. In general, a larger auxiliary sample helps estimate $(\hat{s}_{L},\hat{s}_{U})$ closer to $(s^{*}_{L},s^{*}_{U})$, which improves the length of the target bounds $[\max_{t}\Delta(t,\hat{s}_{L}),1+\min_{t}\Delta(t,\hat{s}_{U})]$. On the other hand, a larger auxiliary sample implies a smaller main sample, which leads to larger variance of $(\widetilde{\theta}_{L},\widetilde{\theta}_{U})$ around the target bounds.
Theorem 2 (Finite-sample validity of sample-splitting CI).
Let Assumption 1 hold. Then, for any $\alpha\in(0,1)$ and any estimators $\hat{s}_{L}$ and $\hat{s}_{U}$, it holds that
$P\left(\theta\in\left[\widetilde{\theta}_{L}-c_{\alpha},1\right]\right)\geq 1-\alpha,\;\;P\left(\theta\in\left[0,\widetilde{\theta}_{U}+c_{\alpha}\right]\right)\geq 1-\alpha,$
$P\left(\theta\in\left[\widetilde{\theta}_{L}-c_{\alpha/2},\widetilde{\theta}_{U}+c_{\alpha/2}\right]\right)\geq 1-\alpha.$
where
$c_{\alpha}=\left(\frac{\log(2/\alpha)}{2}\right)^{\frac{1}{2}}\left(\sqrt{|\mathcal{I}^{1}_{M}|}^{-1}+\sqrt{|\mathcal{I}^{0}_{M}|}^{-1}\right).$
Theorem 2 provides confidence intervals that covers $\theta$ with probability at least $1-\alpha$ for any size of the main and auxiliary samples. Since $\theta\in[0,1]$, the one-sided CIs are particularly useful for testing $\theta=0$ or $\theta=1$. The proof of the theorem exploits sample splitting to take $\hat{s}_{L},\hat{s}_{U}$ as given in the main sample. Then, since by (2.3) $\theta$ must lie within $[\max_{t}\Delta(t,\hat{s}_{L}),1+\min_{t}\Delta(t,\hat{s}_{U})]$, the CIs are built to cover this interval.
The finite-sample guarantee comes from the DKW inequality (Dvoretzky et al., 1956; Massart, 1990), a concentration inequality for empirical cdfs. A similar approach was implemented in Ruiz and Padilla (2022) in the context of finite-sample inference on $\theta$ without covariates. My proof differs in that I consider the splitting of the sample, and I use the one-sided version of the DKW inequality in Massart (1990) to get a smaller critical value for the one-sided CIs.
Note that a large main sample leads to $c_{\alpha}$ being close to zero, and a large auxiliary sample makes a suitable choice of $(\hat{s}_{L},\hat{s}_{U})$ closer to $(s^{*}_{L},s^{*}_{U})$. Therefore, under regularity conditions, the CIs in Theorem 2 collapse into the sharp bounds in (2.5) as both sample sizes grow to infinity. The next section investigates the large sample behavior of a similar estimator, which allows for cross-fitting instead of sample-splitting.
4 Large Sample Inference with Cross-Fitting
The approach of the previous section uses sample-splitting to provide a strong guarantee of coverage under weak conditions. In this section, I propose a new approach that uses cross-fitting to improve power. Unlike sample-splitting, cross-fitting uses the whole sample for estimating both the covariate-adjustment term and the final estimator.
In order to derive asymptotically valid confidence intervals, I define new regularity conditions (A2) and show in Theorem 3 an intermediate result, that the estimators are asymptotically Gaussian. The main result is then established in Theorem 4, which shows uniform validity of CIs for $\theta$. Finally, Theorem 5 shows conditions under which the CIs are exact.
Definition 2 (Cross-fitting estimators).
  1. (i)
    Randomly split the treatment sample into $K$ sets of the same size, that is, a partition of $\{1\leq i\leq n:D_{i}=1\}$ into equal-sized folds $\{\mathcal{I}^{1}_{1},\dots,\mathcal{I}^{1}_{K}\}$. Similarly, split the control sample into $\{\mathcal{I}^{0}_{1},\dots,\mathcal{I}^{0}_{K}\}$, and denote $\mathcal{I}_{k}=\mathcal{I}^{1}_{k}\cup\mathcal{I}^{0}_{k}$;
  2. (ii)
    For all $k\in\{1,\dots,K\}$, use data from all folds except $\mathcal{I}_{k}$ to fit estimators $\hat{s}_{L},\hat{s}_{U}$ to target $s^{*}_{L},s^{*}_{U}$, as in Definition 1.ii. Denote by $\hat{s}_{L,k},\hat{s}_{U,k}$ the estimated functions for each $k$;
  3. (iii)
    Denote by $k(i)$ the index of the fold such that $i\in\mathcal{I}_{k(i)}$, $n_{1}=\sum_{i=1}^{n}D_{i}$ and $n_{0}=\sum_{i=1}^{n}1-D_{i}$. Compute the final estimators
    $\widehat{\theta}_{L}$ $=\max_{t}\Biggl{\{}\frac{1}{n_{1}}\sum_{i:D_{i}=1}\mathbb{I}(Y_{i}\leq\hat{s}_{L,k(i)}(X_{i})+t)-\frac{1}{n_{0}}\sum_{i:D_{i}=0}\mathbb{I}(Y_{i}\leq\hat{s}_{L,k(i)}(X_{i})+t)\Biggr{\}},$
    $\widehat{\theta}_{U}$ $=1+\min_{t}\Biggl{\{}\frac{1}{n_{1}}\sum_{i:D_{i}=1}\mathbb{I}(Y_{i}\leq\hat{s}_{U,k(i)}(X_{i})+t)-\frac{1}{n_{0}}\sum_{i:D_{i}=0}\mathbb{I}(Y_{i}\leq\hat{s}_{U,k(i)}(X_{i})+t)\Biggr{\}}.$
Again, several methods can be used to estimate $(\hat{s}_{L,k},\hat{s}_{U,k})$, including machine learning algorithms or other nonparametric approaches (see, e.g., Kneib et al., 2023 for a review). $K$ is assumed fixed as $n\to\infty$, and typical choices are $K=5$ or $K=10$.
In order to derive the asymptotic properties of the cross-fitting estimators, I define new regularity conditions. The results are uniform over a set of probability functions $\mathcal{P}$. This is because the literature on inference in partially identified models argues that asymptotic results have to be uniform in order to obtain approximations that accurately represent the finite-sample behavior (e.g., Imbens and Manski, 2004; Andrews and Soares, 2010). To this end, I introduce a slight revision in notation, incorporating an additional subscript $P$ for $P\in\mathcal{P}$. Consequently, $\Delta$ in (2.4) is denoted as $\Delta_{P}$, while the sharp transformations in (2.6) and (2.7) are represented by $s^{*}_{L,P}$ and $s^{*}_{U,P}$, and sharp bounds in (2.5) are $(\theta_{L,P}^{*},\theta_{U,P}^{*})$.
Assumption 2.
The set of probability functions $\mathcal{P}$ satisfies:
  1. (i)
    Continuity of distributions of potential outcomes: the sets of cdfs $\{t\mapsto P(Y(j)\leq t):P\in\mathcal{P}\}$ for $j=0,1$ are equicontinuous;
  2. (ii)
    Limit of $\hat{s}_{L}$ and $\hat{s}_{U}$: There exists $s_{L,P},s_{U,P}$ for all $P\in\mathcal{P}$ such that, for all $\varepsilon>0$, $\sup_{P\in\mathcal{P}}P(|\hat{s}_{L}(X)-s_{L,P}(X)|>\varepsilon)\to 0$ and $\sup_{P\in\mathcal{P}}P(|\hat{s}_{U}(X)-s_{U,P}(X)|>\varepsilon)\to 0$ as $n\to\infty$;
  3. (iii)
    Uniqueness of optimizers: There exists $t_{max,P},t_{min,P}$ for all $P\in\mathcal{P}$ such that, for all $\varepsilon>0$,
    $\sup_{P\in\mathcal{P}}\sup_{t:|t-t_{max,P}|\geq\varepsilon}$ $\Delta_{P}(t,s_{L,P})-\Delta_{P}(t_{max,P},s_{L,P})<0,$
    $\sup_{P\in\mathcal{P}}\sup_{t:|t-t_{min,P}|\geq\varepsilon}$ $\Delta_{P}(t,s_{U,P})-\Delta_{P}(t_{min,P},s_{U,P})>0;$
  4. (iv)
    Lower bounded variance: For some $C>0$, $C<P(Y\leq s_{L,P}(X)+t_{max,P})<1-C$ and $C<P(Y\leq s_{U,P}(X)+t_{min,P})<1-C$ for all $P\in\mathcal{P}$.
A2(i) restricts the possibility of point masses in the distribution of the outcome. A2(ii) requires the estimators $(\hat{s}_{L},\hat{s}_{U})$ to have any limit in probability, at any rate of convergence, and allows them to be inconsistent for the sharp $(s^{*}_{L,P},s^{*}_{U,P})$. A2(iii) is assumed to simplify exposition. It ensures the asymptotic distribution is first-order insensitive to the choice of the optimizers in Definition 2(iii), and it is a generalization for $P\in\mathcal{P}$ of the assumption used in Fan and Park (2010) in their approach to inference on Makarov bounds without covariates. I discuss two approaches for relaxing A2(iii) in the presence of multiple optimizers in the Online Appendix (OA.3). One of them allows dropping A2(iii) by using the bootstrap for inference instead of relying on normal approximations to $\widehat{\theta}_{L}$ and $\widehat{\theta}_{U}$. Example 1 illustrates a data generating process where A2(iii) holds. A2(iv) is used to ensure that the problem is not trivial in the sense that $Var[\sqrt{n}\widehat{\theta}_{L}]$ and $Var[\sqrt{n}\widehat{\theta}_{U}]$ do not converge to zero. It is only used to apply the result from Stoye (2009) to get asymptotic validity of the two-sided CI in Theorem 4.
Example 1 (Uniqueness of optimizers).
Consider a data generating process where $Y(0)=f(X)+\varepsilon$, $Y(1)=Y(0)+\tau$, $\varepsilon,\tau\perp\!\!\!\perp X$ and $\mathbb{E}[Y(1)]=\mathbb{E}[Y(0)]=0$. For simplicity, I focus on a single probability function $P$, that is, $\mathcal{P}$ is a singleton. When the distributions of $Y(1)$ and $Y(0)$ are continuous and symmetric around zero, the optimal covariate-adjustment term is the same for both the lower and upper bounds and is given by $s^{*}(x)=\mathbb{E}[Y(0)|X=x]=f(x)$. The distributions of $Y(1)-s^{*}(X)$ (in blue) and $Y(0)-s^{*}(X)$ (in red) are shown in Figure 2 below. The difference $\Delta_{P}(t,s^{*})=P(Y(1)-s^{*}(X)\leq t)-P(Y(0)-s^{*}(X)\leq t)$ is uniquely maximized and minimized respectively at the green and yellow vertical dashed lines, where the probability density functions intersect.
Visualization of Example 1. Blue lines represent the distribution of $Y(1)-s^{*}(X)$, and red lines the distribution of $Y(0)-s^{*}(X)$. The left panel shows the cumulative distribution function and the right panel shows the probability density function. The green dashed line represents the unique point $t$ where $\Delta_{P}(t,s^{*})$ is maximized, and the solid green line in the left panel represents the magnitude of $\Delta_{P}(t,s^{*})$ at this point. Similarly, the yellow dashed and solid lines show the unique point that minimizes $\Delta_{P}(t,s^{*})$ and its magnitude.
Figure 2: Example of data generating process that satisfies A2(iii).
Under A2(iii), the presence of the optimizers in Definition 2(iii) is asymptotically irrelevant, and $(\widehat{\theta}_{L},\widehat{\theta}_{U})$ behave as sample averages. It follows that their asymptotic distribution will be $\sqrt{n}$-Gaussian, as shown in Theorem 3. The presence of the estimated $(\hat{s}_{L,k},\hat{s}_{U,k})$ will only affect where $(\widehat{\theta}_{L},\widehat{\theta}_{U})$ are centered at, but these target bounds will always contain $\theta$, since any $s$ induces valid bounds (2.1).
Theorem 3 (Asymptotic Distribution of Cross-fitting Estimators).
Let $\mathcal{P}$ denote a set of probability functions that satisfy Assumptions 1 and 2. Then, there exists $\bar{\theta}_{L,P},\bar{\theta}_{U,P}$ with $\bar{\theta}_{L,P}\leq\theta\leq\bar{\theta}_{U,P}$ such that
[…]
uniformly in $P\in\mathcal{P}$.
Definitions of $\sigma^{2}_{L,P}$, $\sigma^{2}_{U,P}$ and $\sigma_{L,U,P}$ are provided in Appendix A.5. The cross-fitting estimators are asymptotically normal at the usual $\sqrt{n}$ rate, centered at bounds $[\bar{\theta}_{L,P},\bar{\theta}_{U,P}]$ that may be wider than the sharp $[\theta_{L}^{*},\theta_{U}^{*}]$, but which always contain $\theta$. The nonparametric rates of convergence of $\hat{s}_{L},\hat{s}_{U}$ only affect the target bounds $[\bar{\theta}_{L,P},\bar{\theta}_{U,P}]$, which always contain $\theta$. The resulting $\sqrt{n}$ rate is made possible through the splitting of the sample. Once conditioning on the estimated $(\hat{s}_{L,k},\hat{s}_{U,k})$ and under A2(iii), $(\widehat{\theta}_{L},\widehat{\theta}_{U})$ behave as sample averages. Conditions under which $(\bar{\theta}_{L,P},\bar{\theta}_{U,P})=(\theta_{L,P}^{*},\theta_{U,P}^{*})$ are given in Theorem 5.
Confidence intervals for $\theta$ follow from consistent sample-analogue estimators for $\sigma_{L}^{2}$, $\sigma_{U}^{2}$ and $\sigma_{L,U}$. I provide expressions in Appendix A.5. One-sided confidence intervals follow directly from the normal approximation of Theorem 3. Since $\theta\in[0,1]$, the one-sided CIs are particularly useful for testing $\theta=0$ or $\theta=1$. For two-sided inference on $\theta$, I suggest using the confidence interval $\text{CI}^{3}_{\alpha}$ proposed in Stoye (2009), here denoted by $\widehat{CI}_{\alpha}$. A precise definition is reproduced in Online Appendix OA.2.
Theorem 4 (Uniform Validity of Confidence Intervals).
Let $\mathcal{P}$ denote a set of probability functions that satisfy Assumptions 1 and 2, and denote by $\Theta(P)=[\theta_{L,P}^{*},\theta_{U,P}^{*}]$ the sharp identified set for $\theta$ when $(Y,D,X)\sim P\in\mathcal{P}$. Then,
$\liminf_{n\to\infty}\inf_{P\in\mathcal{P}}\inf_{\theta\in\Theta(P)}P\left(\theta\geq\widehat{\theta}_{L}-z_{\alpha}\frac{\widehat{\sigma}_{L}}{\sqrt{n}}\right)\geq 1-\alpha,$
$\liminf_{n\to\infty}\inf_{P\in\mathcal{P}}\inf_{\theta\in\Theta(P)}P\left(\theta\leq\widehat{\theta}_{U}+z_{\alpha}\frac{\widehat{\sigma}_{U}}{\sqrt{n}}\right)\geq 1-\alpha,$
$\liminf_{n\to\infty}\inf_{P\in\mathcal{P}}\inf_{\theta\in\Theta(P)}P\left(\theta\in\widehat{CI}_{\alpha}\right)\geq 1-\alpha,$
where $z_{\alpha}=\Phi^{-1}(1-\alpha)$, the $(1-\alpha)$-th quantile of the standard normal distribution.
Finally, I show that the target bounds $(\bar{\theta}_{L,P},\bar{\theta}_{U,P})$ converge to the sharp $(\theta^{*}_{L,P},\theta^{*}_{U,P})$ if the estimators of the covariate adjustment terms are consistent to the optimal at any rate. Moreover, I show that the confidence intervals are asymptotically exact under a smoothness and rate condition.
Theorem 5 (Asymptotic exactness of confidence intervals).
[…]
  • (i)
    The cdfs $P(Y(j)\leq t|X=x)$ for $j=0,1$ are twice continuously differentiable in $t$ with second derivative bounded uniformly in $P\in\mathcal{P}$ and $x\in\mathcal{X}$;
  • (ii)
[…]
Note that a simple sufficient condition for (ii) in Theorem 5 is that $|\hat{s}_{L,k}(X)-s^{*}_{L,P}(X)|$ converges to zero at the semiparametric rate $o_{P}(n^{-1/4})$.
5 Monte Carlo Simulations
I illustrate the results of Theorems 2 and 3 with a simulation study. I also compare the performance of the sample-splitting and cross-fitting estimators, as well as the estimator proposed by Semenova (2023) and Ji et al. (2023) (denoted SJLS for the first letter of the authors’ names). Note that both estimators are equivalent in the case of the DTE and that CAIDE by construction estimates bounds that are always (weakly) narrower than SJLS (Appendix A.1).
For $n\in\{500,2000\}$, I simulate $10,000$ times from the following data-generating process (DGP):
  • $d=20$ covariates:
    • $\circ$
      $X\sim\mathcal{N}(0,\Sigma)$, $\Sigma_{ij}=0$ if one of $i$ or $j\in\{1,2\}$, $\Sigma_{ij}=0.5^{|i-j|}$ otherwise.
  • Potential outcomes:
    • $\circ$
      $\beta_{0}=(3^{1},3^{0},0,0,\dots,0,3^{-1},3^{-2},\dots,3^{-6})$, $\Theta_{0,ij}=0.2(-1)^{i+j}$;
    • $\circ$
      $\beta_{\tau}=(1,1,\dots,1)$, $\Theta_{\tau,ij}=0.2$;
    • $\circ$
      $Y_{i}(0)=X^{\prime}_{i}\beta_{0}+X^{\prime}_{i}\Theta_{0}X_{i}$;
    • $\circ$
      $Y_{i}(1)=Y_{i}(0)-1+X^{\prime}_{i}\beta_{\tau}+X^{\prime}_{i}\Theta_{\tau}X_{i}$;
    • $\circ$
      $Y_{i}=D_{i}Y_{i}(1)+(1-D_{i})Y_{i}(0)$, $D_{i}\sim\text{Bernoulli}(0.5)$.
For this DGP, $\theta=\theta_{0}\approx 0.43$. I fit with $K=5$ and $\delta=0$ the models Quantile Neural Networks, Support Vector Machine (Regression), and No Covariates ($s(x)=0$). I also calculate an Oracle model that uses the sharp transformations defined in Theorem 1, which are generally unknown if the DGP is unknown. Technical details are delayed to Appendix A.4. I consider two scenarios: (i) $p=10$, where only covariates $X_{1}$ through $X_{10}$ are observed; (ii) $p=20$, where all $20$ covariates are observed. Note that when $p=20$, $\theta$ is point identified since potential outcomes are uniquely determined by all $20$ covariates.
Note: Green dashed lines represent the sharp identified set of $\theta$ when $p=10$, and orange dashed lines represent the true value of $\theta$. Boxplots comprise the median, two hinges at the first and third quartiles, and whiskers extending to the most extreme data points no further than 1.5 times the interquartile range (third minus first quartile) from the hinge.
Figure 3: Distribution of Estimators for Lower and Upper Bounds on $\theta$
Figure 3 shows the distribution of the estimators CAIDE-CF (Cross Fitting) and SJLS for the lower and upper bounds on $\theta$, for $n=500,2000$ and scenarios $p=10,20$. It illustrates that the estimators are centered around bounds determined by the different models for estimating $\hat{s}$. Some models perform better than others in yielding narrower bounds. For example, in most cases, the Regr. SVM model led to the narrowest bounds, close to the sharp bounds in the $p=10$ case. A larger sample size leads to smaller variation in the estimators. Observing all $p=20$ covariates, instead of $p=10$, can lead to narrower bounds (e.g., Regr. SVM and $n=2000$) but also to wider bounds if the sample is smaller (Quant. NNet and $n=500$). Throughout, the estimates of CAIDE are narrower than those of SJLS.
In Table 1, I show the probability of rejecting the hypotheses $H_{0}:\theta=0$ and $H_{0}:\theta\geq\theta_{0}$, using the one-sided confidence interval based on the lower bound in Theorem 4. Here, $\theta_{0}\approx 0.43$ is the true value of $\theta$ for this DGP. A high rejection rate for the first hypothesis indicates power since 00 is not in the sharp identified set of $\theta$. The second hypothesis’s rejection rate indicates size, and it should ideally be less than or equal to the nominal $0.05$. I also measure the average length of the confidence intervals, defined as the distance between the lower and upper bounds calculated from the one-sided CIs in Theorem 4. I compare $n=500,2000$ and $p=10,20$ for different models and estimators. CAIDE-SS is the sample-splitting estimator of Section 3, and CAIDE-CF is the cross-fitting estimator of Section 4. The table shows that including covariates, having a larger sample size, or observing all $20$ covariates increases power while preserving size. The sample-splitting estimators are conservative and exhibit the largest average length for Regr. SVM. Still, it often demonstrates some power. For example, it excludes zero in $99.4\%$ of the cases when $p=20$, $n=2000$, and the model is Regr. SVM. The cross-fitting estimator performs better than SJLS in terms of smaller confidence intervals while preserving coverage. It achieves the nominal coverage rate for the oracle model when $p=20$.
Table 1: Measures of power and coverage for lower bound on $\theta$
p = 10 p = 20
n Estimator $\theta=0$ $\theta\geq\theta_{0}$ Avg Length $\theta=0$ $\theta\geq\theta_{0}$ Avg Length
No Covariates
$500$ Sample Split. $0.000$ $0.000$ $0.998$ $0.000$ $0.000$ $0.998$
$500$ Cross Fit. $0.722$ $0.000$ $0.802$ $0.722$ $0.000$ $0.802$
$500$ SJLS $0.000$ $0.000$ $1.000$ $0.000$ $0.000$ $1.000$
$2,000$ Sample Split. $0.001$ $0.000$ $0.892$ $0.001$ $0.000$ $0.892$
$2,000$ Cross Fit. $0.983$ $0.000$ $0.776$ $0.983$ $0.000$ $0.776$
$2,000$ SJLS $0.000$ $0.000$ $0.965$ $0.000$ $0.000$ $0.965$
Regr. SVM
$500$ Sample Split. $0.004$ $0.000$ $0.935$ $0.007$ $0.000$ $0.948$
$500$ Cross Fit. $0.970$ $0.000$ $0.698$ $0.998$ $0.000$ $0.676$
$500$ SJLS $0.806$ $0.000$ $0.815$ $0.964$ $0.000$ $0.791$
$2,000$ Sample Split. $0.414$ $0.000$ $0.804$ $0.994$ $0.000$ $0.716$
$2,000$ Cross Fit. $1.000$ $0.000$ $0.628$ $1.000$ $0.000$ $0.485$
$2,000$ SJLS $1.000$ $0.000$ $0.707$ $1.000$ $0.000$ $0.562$
Oracle
$500$ Cross Fit. $0.999$ $0.000$ $0.608$ $1.000$ $0.051$ $0.103$
$2,000$ Cross Fit. $1.000$ $0.000$ $0.568$ $1.000$ $0.049$ $0.052$
6 Application to Microcredit
In 2022, microcredit reached more than 170 million borrowers worldwide, with a total loan portfolio of over \$140 billion (Convergences, 2023). Theory suggests that microborrowing can be beneficial or harmful (Banerjee, 2013, Garz et al., 2021), and evidence suggests that average effects are small if not negligible (Banerjee et al., 2015b, Meager, 2019). Analysis of heterogeneity mostly relied on rank preservation and suggested mainly positive impact in the upper tail of income/profits distributions (e.g., Meager, 2022). Yet, evidence of negative effects has been scarce. I revisit five studies that randomized offering a microloan at the individual level [^5] The only exception is Egypt, in which the treated group was offered a loan size twice as large as the one offered in the control group., and find evidence of both positive and negative treatment effects of microcredit. Table 2 presents a brief description of the datasets.
CAIDE is able to provide informative bounds, even in this challenging context. First, the setting is challenging because the distributions of income/profits are similar between the treated and control groups. This similarity is reflected in not statistically significant ATEs in Table 2, and estimated $[\theta_{L},\theta_{U}]$ close to $[0,1]$ if covariates are not considered (Table 3). Second, because the sample sizes are relatively small (ranging from 548 to 1964), with a large number of covariates (from 24 to 97). Finally, because predicting business outcomes is itself a difficult task (McKenzie and Sansone, 2019).
Table 2: Description of datasets in microcredit
Country Bosnia Egypt Philippines (1) Phil. (2) S. Africa
Year of implementation 2009 2016 2006 2010 2004
Treated/control obs. 551/444 477/478 891/222 1651/313 235/313
Number of covariates 48 97 24 39 33
Avg. loan size in treatment group (USD PPP) \$1,389 \$9,912 \$1,045 \$492 \$180
Avg household income at endline (USD PPP) \$21,735 \$26,844 \$12,173 \$10,280 \$14,845
ATE in Household Income (t-statistic)
\$609
\$1,224
\$1,374
\$110
-\$520
(-0.159)
  • Study citations respectively: Augsburg et al. (2015), Bryan et al. (2021), Karlan and Zinman (2011), Karlan et al. (2016), Karlan and Zinman (2010).
Using $K=10$, I fit the estimators of Section 4666The sample-splitting estimators of Section 3 with a 50-50 split lead all confidence intervals to be $[0,1]$ at $10\%$ significance level.. For each fold, I fit another layer of sample splitting to pick via $10-$fold cross-validation one of the models: Quantile Forests, Quantile Neural Networks, Random Forest (Regression), Neural Networks (Reg.), Elastic Net (Reg.), Extreme Gradient Boosting (Reg.), and Constant ($s(X)=0)$. Technical details are provided in Appendix A.4.
Table 3: Bounds on fraction harmed by microcredit - household income
Lower Bound Upper Bound
CAIDE No Covariates SJLS CAIDE No Covariates SJLS
Bosnia
0.108
[0.000]
0.040
[0.073]
0.024
[0.254]
0.946
[0.028]
0.973
[0.195]
0.977
[0.310]
Egypt
0.260
[0.000]
0.009
[0.263]
0.248
[0.000]
0.787
[0.000]
0.966
[0.051]
0.792
[0.000]
Philippines (1)
0.083
[0.018]
0.071
[0.036]
0.077
[0.075]
0.971
[0.057]
0.931
[0.005]
1.048
[1.000]
Philippines (2)
0.145
[0.000]
0.022
[0.103]
0.075
[0.013]
0.739
[0.000]
0.951
[0.062]
0.751
[0.000]
South Africa
0.028
[0.053]
0.033
[0.036]
-0.017
[1.000]
0.876
[0.002]
0.913
[0.021]
0.938
[0.105]
All
0.125
[0.000]
0.035
[0.000]
0.081
[0.000]
0.864
[0.000]
0.947
[0.000]
0.901
[0.000]
  • Standard errors in parentheses, and p-values in brackets. The null hypothesis for the lower bound is $\theta^{*}_{L}=0$, and for the upper bound $\theta^{*}_{U}=1$. Response variable: $log(1+household\ income)$.
Table 3 shows the estimates for the lower and upper bounds on $\theta$ for $\delta=0$ using CAIDE with and without covariates [^7] Note that the estimators and one-sided CIs without covariates ($s(x)=0$) are equivalent to the ones in Fan and Park (2010).. The outcome is the logarithm of annual household income. Considering that the outcome is continuous, it follows that $\theta=P(Y(1)-Y(0)\leq 0)=P(Y(1)-Y(0)<0)$, the lower bound on $\theta$ is the minimum proportion of households that have a negative effect from the microloan, and the upper bound is the maximum proportion. Note that one minus the upper bound denotes the minimum proportion of households with a positive effect. Results for $\delta=-0.05$ and $\delta=0.05$ are similar and are presented in Table 4. In Bosnia, for example, I estimate that at least $10.8\%$ of the households would be better off without the microloan (p-value $<10^{-3}$), and at most $94.6\%$ have negative effects (p-value $0.028$). That is, at least $6.4\%$ are better off ($1-0.946$). Overall, I find significant lower bounds at the $10\%$ level for all five datasets and at the $5\%$ level for four of them. As for the upper bound, I find statistical significance at $5\%$ in four datasets. These results reveal an important presence of heterogeneity, even when the ATEs are not statistically significant (Table 2).
Incorporating covariates makes a substantial difference in many of the point estimates. In Egypt, for example, the lower bound without covariates is $0.009$ (p-value $0.264$), whereas incorporating covariates improves the estimate to $0.260$ (p-value $<10^{-3}$). The differences between the point estimates with and without covariates are significant at the $5\%$ level for both the lower and upper bounds for Egypt, the Philippines (2) and All.
Table 3 also illustrates how CAIDE can outperform SJLS (Semenova, 2023, Ji et al., 2023) in terms of narrower confidence intervals. In the case of the lower bound for Bosnia, for example, it is estimated $10.8\%$ with CAIDE vs. $2.4\%$ with SJLS, leading to a substantial difference in p-values ($<10^{-3}$ vs. $0.254$). It also illustrates how SJLS can estimate negative lower bounds (South Africa), or upper bounds greater than one (Philippines (1)). In both cases, the estimates with CAIDE are statistically significant at the $10\%$ level.
Table 5 in Online Appendix OA.4 shows results for business profits. Again, including covariates improves the bounds. Averaging estimates from all datasets, I find that at least $10.8\%$ (p-value $<10^{-3}$) of the population was harmed, and at least $9.8\%$ (p-value $<10^{-3}$) benefited from microcredit.
An important note is that this evidence does not inform the nature of the negative effects or their magnitude. For example, the negative outcomes could be relatively small and a natural consequence of investments being risky. In that case, the decision to take a microloan could still be rational if the prospect of gains is better than the risk of loss.
7 Conclusion
I propose two novel inference approaches to the distributional treatment effect $\theta=P(Y(1)-Y(0)\leq\delta)$ for any fixed $\delta$. A new characterization of the sharp identified set of $\theta$ is provided, based on covariate adjustment. It has the desirable property of being robust to the choice of the covariate-adjustment term $s(x)$. If this term is misspecified, the identified set is wider but still valid. As a consequence, any method can be used to estimate $s$, including machine learning algorithms. I provide finite-sample valid inference using sample-splitting and asymptotic inference using cross-fitting, which is also asymptotically exact under additional conditions.
I illustrate the practical relevance of the method in a simulation study and an application to microcredit. The application reveals an important presence of heterogeneity, with statistically significant negative and positive treatment effects. The results show that the method is able to provide informative bounds even in challenging contexts, such as when the ATE is not statistically significant, the sample size is small, and the number of covariates is large.
By using CAIDE, an applied researcher can learn more about heterogeneous effects not only by gathering larger samples but also by collecting information on variables that help predict potential outcomes.
Appendix A
A.1 Comparison with Semenova (2023) and Ji et al. (2023)
I compare the asymptotic properties of CAIDE with the estimators proposed by Semenova (2023) and Ji et al. (2023) (denoted by SJLS for the first letter of the names of the authors). I focus on the cross-fitting version of my estimators (Definition 2(iii)) and compare them to the cross-fitting estimators in SJLS. Note that in the case of the DTE, $\theta$, the estimators of Semenova (2023) and Ji et al. (2023) are equivalent.
Since SJLS considered a known propensity score, I compare SJLS to a version of CAIDE that uses the true propensity score (Online Appendix OA.1) instead of using the fraction of treated units in the sample. In Appendix A.2, I show that using the actual number of treated units (as in Definition 2(iii)) instead of the known propensity score leads to a smaller asymptotic variance. Hence, using the true propensity score in CAIDE is the natural comparison. I focus on the lower bound for simplicity, and analogous results hold for the upper bound. Define
$\widehat{\Delta}(t,\hat{s}_{L})=\frac{1}{n}\sum_{i=1}^{n}\biggl{(}\frac{D_{i}}{p(X_{i})}\mathbb{I}(Y_{i}\leq\hat{s}_{L,k(i)}(X_{i})+t)-\frac{1-D_{i}}{1-p(X_{i})}\mathbb{I}(Y_{i}\leq\hat{s}_{L,k(i)}(X_{i})+t)\biggr{)},$
where $\hat{s}_{L}=(\hat{s}_{L,1},\dots,\hat{s}_{L,K})$.
Then, CAIDE is equivalent to $\hat{\theta}_{C}:=\max_{t}\widehat{\Delta}(t,\hat{s}_{L})$ and SJLS is given by $\hat{\theta}_{S}:=\widehat{\Delta}(0,\hat{s}_{L})$. Theorem 3 of this paper is comparable to Theorem 4.1 in Semenova (2023), and Theorem 3.1 and Proposition 3.2 in Ji et al. (2023). They all show asymptotic validity of confidence intervals. Define confidence intervals as $CI_{C}$ for CAIDE and $CI_{S}$ for SJLS. Since by construction $\hat{\theta}_{C}\geq\hat{\theta}_{S}$, it follows that $\hat{\theta}_{C}$ first order stochastically dominates $\hat{\theta}_{S}$. Let $\theta_{n}^{*}=\Delta_{P_{n}}(0,s^{*}_{L})$ be the sharp lower bound as in (2.5) and Theorem 1. Under the conditions of these theorems, it follows that, for any sequence $\theta_{n}<\theta_{n}^{*}$,
$\liminf_{n\to\infty}P_{n}(\theta_{n}\in CI_{S})-P_{n}(\theta_{n}\in CI_{C})\geq 0.$
Furthermore, consider a scenario where $\max_{t}\Delta_{P_{n}}(t,s_{L,P})>\Delta_{P_{n}}(0,s_{L,P})+M$ for some $M>0$, where $s_{L,P}$ is the limit of $\hat{s}_{L,k}$ (as in A2(ii)). This is usually the case when $s_{L,P}$ is not a sharp transformation $s^{*}_{L,P}$. Then, for any $\Delta_{P_{n}}(0,s_{L,P_{n}})+M\leq\theta_{n}\leq\max_{t}\Delta_{P_{n}}(t,s_{L,P_{n}})$, the inequality above holds with strict inequality.
Since $\theta_{n}<\theta_{n}^{*}$, a lower coverage probability translates into higher power. Hence, $CI_{C}$ dominates $CI_{S}$ in terms of power. Note that validity of $CI_{C}$ relies on A2(iii), on the uniqueness of the maximizer of $\Delta_{P_{n}}(t,s_{L,P})$, which is not assumed in SJLS. I discuss two approaches to weaken this assumption in Online Appendix OA.3.
A.2 Estimated vs True Propensity Score
I show that when the propensity score is constant, the asymptotic variances of the estimators in 2(iii) are smaller than those of alternative estimators that consider the true propensity score. When the propensity score is known and not constant as in A1.ii, an alternative estimator that uses the true propensity score is proposed in Online Appendix OA.1-B. The lower bound, for example, is given by
$\widehat{\theta}^{B}_{L}$ $=\max_{t}\Biggl{\{}\frac{1}{n}\sum_{i=1}^{n}\biggl{(}\frac{D_{i}}{p(X_{i})}\mathbb{I}(Y_{i}\leq\hat{s}_{L,k(i)}(X_{i})+t)$
$\hskip 60.0pt-\frac{1-D_{i}}{1-p(X_{i})}\mathbb{I}(Y_{i}\leq\hat{s}_{L,k(i)}(X_{i})+t)\biggr{)}\Biggr{\}}.$
Now, note that the estimators in Definition 2(iii) can be rewritten in terms of an estimated propensity score $\hat{p}=\frac{1}{n}\sum_{i=1}^{n}D_{i}$. For the lower bound, we have
$\widehat{\theta}_{L}$ $\equiv\max_{t}\Biggl{\{}\frac{1}{n}\sum_{i=1}^{n}\biggl{(}\frac{D_{i}}{\hat{p}}\mathbb{I}(Y_{i}\leq\hat{s}_{L,k(i)}(X_{i})+t)$
$\hskip 60.0pt-\frac{1-D_{i}}{1-\hat{p}}\mathbb{I}(Y_{i}\leq\hat{s}_{L,k(i)}(X_{i})+t)\biggr{)}\Biggr{\}}.$
Now, let $t_{L,P}=t_{max,P}$; $t_{U,P}=t_{min,P}$; $\pi=P(D=1)$ (as in A1). For $A\in\{L,U\}$ and $j\in\{0,1\}$, define $Z_{A,P}(j)=\mathbb{I}(Y(j)\leq s_{A,P}(X)+t_{A,P})$. Then, the asymptotic variance of $\widehat{\theta}_{A}$ (as in Definition 2(iii)) is given by:
$\frac{Var[Z_{A,P}(1)]}{\pi}+\frac{Var[Z_{A,P}(0)]}{1-\pi}$
Now, if $p(x)=\pi$, the asymptotic variance of $\widehat{\theta}^{B}_{A}$ simplifies to:
$\frac{Var[Z_{A,P}(1)]}{\pi}+\frac{Var[Z_{A,P}(0)]}{1-\pi}+\left(\frac{\mathbb{E}[Z_{A,P}(1)]}{\pi}-\frac{\mathbb{E}[Z_{A,P}(0)]}{1-\pi}\right)^{2}\pi(1-\pi)$
Since the third term above is never negative and, in general, different from zero, the asymptotic variance of $\widehat{\theta}_{A}$ is smaller than the one of $\widehat{\theta}^{B}_{A}$.
A.3 Asymptotic Distribution of Plug-in Estimator Based on Fan and Park (2010)
I argue that it is generally difficult to characterize the asymptotic distribution of a plug-in estimator based on Fan and Park (2010)’s representation of the identified set (2.5) when the conditional distributions are estimated nonparametrically. I show that the asymptotic distribution depends on the rate of convergence of the estimators of $F_{1}(t|x)$ and $F_{0}(t|x)$, and that different models/algorithms used to estimate these functions lead to different asymptotic distributions of the plug-in estimator. This is true even when using sample-splitting. For simplicity, I focus on the lower bound.
Define
$\Delta(t|x)=F_{1}(t|x)-F_{0}(t|x).$
Let $\hat{F}_{1}(t|x)$ and $\hat{F}_{0}(t|x)$ be estimators respectively of $F_{1}(t|x)$ and $F_{0}(t|x)$, and $\hat{\Delta}(t|x)=\hat{F}_{1}(t|x)-\hat{F}_{0}(t|x)$. Define $\hat{s}(x)\in\arg\min_{t}\hat{\Delta}(t|x)$. Since my estimation strategies in Sections 3 and 4 use sample-splitting, I also use sample-splitting to define the plug-in estimator based on (2.5). Consider a random split of $\{1,\dots,n\}$ into three equal-sized sets $\mathcal{I}_{1}$, $\mathcal{I}_{2}$, and $\mathcal{I}_{3}$ ($n$ is assumed a multiple of 3 for simplicity). Let $\hat{\Delta}$ be estimated using data from $\mathcal{I}_{1}$ only, and $\hat{s}$ estimated using data from $\mathcal{I}_{2}$ only. The plug-in estimator of $\theta_{L}^{*}$ based on (2.5) is given by
$\hat{\theta}_{FP}=\frac{1}{n/3}\sum_{i\in\mathcal{I}_{3}}\hat{\Delta}(\hat{s}(X_{i})|X_{i}).$
Let $\mathcal{D}_{j}=(Y_{i},D_{i},X_{i})_{i\in\mathcal{I}_{j}}$ for $j=1,2,3$. Define
$\bar{\Delta}(\mathcal{D}_{1},\mathcal{D}_{2})=\mathbb{E}[\hat{\Delta}(\hat{s}(X)|X)|\mathcal{D}_{1},\mathcal{D}_{2}],$
where $X$ is independent of $\mathcal{D}_{1},\mathcal{D}_{2}$. It follows that
$\left(\hat{\theta}_{FP}-\theta_{L}^{*}\right)=\left(\frac{1}{n/3}\sum_{i\in\mathcal{I}_{3}}\hat{\Delta}(\hat{s}(X_{i})|X_{i})-\bar{\Delta}(\mathcal{D}_{1},\mathcal{D}_{2})\right)+\left(\bar{\Delta}(\mathcal{D}_{1},\mathcal{D}_{2})-\theta_{L}^{*}\right).$
Since the first term is mean-zero, it follows that
$\sqrt{n}\left(\frac{1}{n/3}\sum_{i\in\mathcal{I}_{3}}\hat{\Delta}(\hat{s}(X_{i})|X_{i})-\bar{\Delta}(\mathcal{D}_{1},\mathcal{D}_{2})\right)$
is asymptotically normal by a triangular array CLT. The second term, $\bar{\Delta}(\mathcal{D}_{1},\mathcal{D}_{2})-\theta_{L}^{*}$, is random since the data $(\mathcal{D}_{1},\mathcal{D}_{2})$ is random. $\bar{\Delta}(\mathcal{D}_{1},\mathcal{D}_{2})-\theta_{L}^{*}$ may or may not be bounded in probability when multiplied by $\sqrt{n}$. In general, the rate of convergence of this term depends on the rate of convergence of $\hat{\Delta}$. Since $\hat{\Delta}$ is estimated nonparametrically, this rate may be slow when regularity conditions such as sparsity and specific forms of smoothness don’t hold. In contrast, the asymptotic distribution in Theorem 3 does not rely on such conditions. The asymptotic distribution of an appropriately scaled version of $\bar{\Delta}(\mathcal{D}_{1},\mathcal{D}_{2})-\theta_{L}^{*}$ depends on which model/algorithm is used to estimate $\hat{\Delta}$, since different estimators $\hat{\Delta}$ lead to different values of $\bar{\Delta}$. For example, the distribution of $\bar{\Delta}(\mathcal{D}_{1},\mathcal{D}_{2})$ is different and difficult to characterize if $\hat{\Delta}$ is calculated with random forests or neural networks, which are some of the models I use in Section 6.
Note that the term $\bar{\Delta}(\mathcal{D}_{1},\mathcal{D}_{2})-\theta_{L}^{*}$ may be positive or negative. If $\bar{\Delta}(\mathcal{D}_{1},\mathcal{D}_{2})$ is symmetrically distributed around $\theta_{L}^{*}$, $\bar{\Delta}(\mathcal{D}_{1},\mathcal{D}_{2})-\theta_{L}^{*}>0$ with probability $1/2$. In this case, a hypothesis test for $H_{0}:\theta_{L}^{*}\leq\tau$, $\tau\in[0,1]$, based on the asymptotic distribution of $\sqrt{n}\left(\frac{1}{n/3}\sum_{i\in\mathcal{I}_{3}}\hat{F}(\hat{s}(X_{i})|X_{i})-\bar{\Delta}(\mathcal{D}_{1},\mathcal{D}_{2})\right)$ may overreject. That is, if one conducts a hypothesis test with significance level $\alpha$ based on the incorrect assumption that $\sqrt{n}\left(\bar{\Delta}(\mathcal{D}_{1},\mathcal{D}_{2})-\theta_{L}^{*}\right)=o_{P}(1)$, the test may, for example, reject $\theta_{L}^{*}=0$ when $\theta=0$ with probability higher than $\alpha$.
In contrast, my approaches to inference do not depend on the rate of convergence or the specific model/algorithm used to estimate $s^{*}_{L}$. Theorem 2 provides a finite-sample guarantee of coverage under no restrictions on how $\hat{s}_{L}$ is calculated. In Theorem 4, I provide a large-sample guarantee of coverage under no restrictions on the rate of convergence of $\hat{s}_{L}$ or which model/algorithm is used to calculate it. The result of Theorem 4 is possible because of the alternative characterization of the identified set in Theorem 1. Unlike (2.5), which relies on the average of conditional expectations, (2.3) is expressed in terms of unconditional expectations. As a consequence, Theorem 3 shows that estimation error in $\hat{s}_{L}$ translates into negative bias,
$\mathbb{E}[\sqrt{n}(\widehat{\theta}_{L}-\theta_{L}^{*})]\leq 0+o_{P}(1),$
which makes the estimate $\widehat{\theta}_{L}$ conservative. Since $\theta_{L}^{*}$ is a lower bound, an estimate $\widehat{\theta}_{L}$ that is too low leads to a confidence set that may be conservative, but valid.
A.4 Estimation of the covariate-adjustment term
I consider two approaches to estimate the covariate-adjustment term $s(x)$. For simplicity, I focus the explanation on the case of the lower bound. First, I estimate $\hat{F}_{j}(t|x)$ ($j=0,1$) with one of two methods, and then I calculate $\hat{s}$ from $\hat{F}_{j}(t|x)$.
The first method for estimating $\hat{F}_{j}(t|x)$ is using quantile models, such as quantile forests or neural networks. They give estimates for quantiles, that is, $\hat{F}^{-1}_{j}(\tau|x)$, for all values of $x$ in the sample. $\hat{F}_{j}(t|x)$ can be calculated from $\hat{F}^{-1}_{j}(\tau|x)$ by linear interpolation. In Sections 5 and 6, I consider a grid for $\tau$ from 00 to $1$ by increments of $0.01$, i.e. $\tau_{k}\in\{0,0.01,\dots,1\}$. For an arbitrary $t$, pick $\tau_{a}$ and $\tau_{b}$ such that $t_{a}=\hat{F}^{-1}_{j}(\tau_{a}|x),t_{b}=\hat{F}^{-1}_{j}(\tau_{b}|x)$ are the two closest values to $t$ such that $t_{a}\leq t\leq t_{b}$. Then, $\hat{F}_{j}(t|x)=\tau_{a}+\frac{\tau_{b}-\tau_{a}}{t_{b}-t_{a}}(t-t_{a})$.
The second method for estimating $\hat{F}_{j}(t|x)$ assumes that the distribution of $Y(j)|X=x$ is the same for all $x$ up to a shift in the mean $\mathbb{E}[Y(j)|X=x]$. This is allowed by the framework of Theorems 2 and 3 since misspecification of $s^{*}$ leads to valid even if wider bounds. Denote by $\hat{\mu}(x)$ an estimator of $\mathbb{E}[Y(j)|X=x]$, using any regression method. For example, in Section 5 I use Support Vector Machine. Denote by $\hat{F}_{e}(t)$ the empirical cdf of the residuals $Y(j)-\hat{\mu}(X)$ inside the training sample. Then, $\hat{F}_{j}(t|x)=\hat{F}_{e}(t-\hat{\mu}(x))$.
Once $\hat{F}_{j}(t|x)$ is estimated, the covariate-adjustment term is calculated by $\hat{s}(x)=\arg\max_{t\in\mathcal{T}}\left\{\hat{F}_{1}(t|x)-\hat{F}_{0}(t|x)\right\}$. For $\mathcal{T}$, I use a random grid of size $10^{4}$ from a normal distribution of mean 00 and standard deviation $\max(Y)-\min(Y)$.
I fit all models in R. For quantile neural networks, I use the package qrnn with 3 hidden nodes, in 1 trial and a maximum of 100 iterations. Other parameters are the default choice. For quantile random forests, I use the package ranger with 1,000 trees and otherwise default hyperparameters. For all the regression methods, I use the package mlr3 with default values of hyperparameters. The packages for all models, used through mlr3 are: random forests (ranger), extreme gradient boosting (xgboost), support vector machine (e1071), neural networks (nnet), elastic net (glmnet). The constant/none model takes $s(x)=0$.
For choosing the best model in the application of Section 6, I use cross-validation inside each fold. That is, for each fold, I fit all the models in a 10-fold cross-validation and choose the model that gives the largest (smallest) lower (upper) bound. Note that the test set is never used in this approach.
When calculating the sample-splitting estimators and confidence intervals of Section 3, I use a 50-50 split of the sample.
A.5 Definitions Ommited from the Main Text
Let $t_{L,P}=t_{max,P}$; $t_{U,P}=t_{min,P}$; $\pi=P(D=1)$ (as in A1); For $A\in\{L,U\}$ and $j\in\{0,1\}$, define $Z_{A,P}(j)=\mathbb{I}(Y(j)\leq s_{A,P}(X)+t_{A,P})$; $\sigma^{2}_{A,P}(j)=Var[Z_{A,P}(j)]$. Then, for $A\in\{L,U\}$:
$\sigma^{2}_{A,P}=\frac{Var[Z_{A,P}(1)]}{\pi}+\frac{Var[Z_{A,P}(0)]}{1-\pi}$
and
$\sigma_{L,U,P}=\frac{Cov[Z_{L,P}(1),Z_{U,P}(1)]}{\pi}+\frac{Cov[Z_{L,P}(0),Z_{U,P}(0)]}{1-\pi}$
For the sample analogues, for $A\in\{L,U\}$ define
$\widehat{F}_{A,j}(t)=\frac{1}{n_{j}}\sum_{i:D_{i}=j}\mathbb{I}(Y_{i}\leq\hat{s}_{A,k(i)}(X_{i})+t);$
$\hat{t}_{L}=\arg\max_{t}\widehat{F}_{A,1}(t)-\widehat{F}_{A,0}(t);\ \hat{t}_{U}=\arg\min_{t}\widehat{F}_{A,1}(t)-\widehat{F}_{A,0}(t);$
$\hat{\sigma}^{2}_{A,j}=\frac{1}{n_{j}}\sum_{i:D_{i}=j}\left[\mathbb{I}(Y_{i}\leq\hat{s}_{A,k(i)}(X_{i})+\hat{t}_{A})-\widehat{F}_{A,j}(\hat{t}_{A})\right]^{2};$
$\hat{\sigma}_{L,U,j}=$
$\frac{1}{n_{j}}\sum_{i:D_{i}=j}\left[\mathbb{I}(Y_{i}\leq\hat{s}_{L,k(i)}(X_{i})+\hat{t}_{L})-\widehat{F}_{L,j}(\hat{t}_{L})\right]\left[\mathbb{I}(Y_{i}\leq\hat{s}_{U,k(i)}(X_{i})+\hat{t}_{U})-\widehat{F}_{U,j}(\hat{t}_{U})\right];$
and $\hat{\pi}=n_{1}/n$. Then, for $A\in\{L,U\}$,
$\hat{\sigma}^{2}_{A}=\frac{\hat{\sigma}^{2}_{A,1}}{\hat{\pi}}+\frac{\hat{\sigma}^{2}_{A,0}}{1-\hat{\pi}};\quad\hat{\sigma}_{L,U}=\frac{\hat{\sigma}_{L,U,1}}{\hat{\pi}}+\frac{\hat{\sigma}_{L,U,0}}{1-\hat{\pi}}.$
A.6 Proofs of the Main Results
Proof of Theorem 2.
For the case of the lower bound, denote for $j=0,1$
$\hat{F}_{j,L}(t)=\frac{1}{|\mathcal{I}^{j}_{M}|}\sum_{i\in\mathcal{I}^{j}_{M}}\mathbb{I}\left(Y_{i}-\hat{s}_{L}(X_{i})\leq t\right)$
Let $\mathcal{I}_{A}$ denote the auxiliary sample, and let $F_{j,L}(t)=\mathbb{E}[\hat{F}_{j,L}(t)|\mathcal{I}_{A}]$ for $j=0,1$ denote the true cdf of $Y(j)-\hat{s}_{L}(X)$ taking $\hat{s}_{L}$ as fixed. Let $\hat{t}_{max}\in\arg\max_{t}[\hat{F}_{1,L}(t)-\hat{F}_{0,L}(t)]$.
Then, we have that $\widetilde{\theta}_{L}=\hat{F}_{1,L}(\hat{t}_{max})-\hat{F}_{0,L}(\hat{t}_{max})$. Finally, denote $c_{\alpha,1}=\left(\frac{\log(2/\alpha)}{2|\mathcal{I}^{M}_{1}|}\right)^{1/2}$, $c_{\alpha,0}=\left(\frac{\log(2/\alpha)}{2|\mathcal{I}^{M}_{0}|}\right)^{1/2}$.
I first focus on the case of the one-sided interval $P\left(\theta\in\left[\widetilde{\theta}_{L}-c_{\alpha},1\right]\right)$. Note that $F_{1,L}(t)-F_{0,L}(t)$ is a valid lower bound by (2.3) for any $t$ and $\hat{s}_{L}$. Thus, $\theta\in[F_{1,L}(\hat{t}_{max})-F_{0,L}(\hat{t}_{max}),1]$ and $F_{1,L}(\hat{t}_{max})-F_{0,L}(\hat{t}_{max})\leq\theta$. Then,
$P\left(\theta\in\left[\widetilde{\theta}_{L}-c_{\alpha},1\right]\right)$
$\geq P\left(F_{1,L}(\hat{t}_{max})-F_{0,L}(\hat{t}_{max})\geq\widetilde{\theta}_{L}-c_{\alpha}\right)$
$=1-P\left(\left[\hat{F}_{1,L}(\hat{t}_{max})-F_{1,L}(\hat{t}_{max})\right]+\left[F_{0,L}(\hat{t}_{max})-\hat{F}_{0,L}(\hat{t}_{max})\right]>c_{\alpha}\right)$
$\geq 1-P\left(\left[\hat{F}_{1,L}(\hat{t}_{max})-F_{1,L}(\hat{t}_{max})\right]>c_{\alpha,1}\lor\left[F_{0,L}(\hat{t}_{max})-\hat{F}_{0,L}(\hat{t}_{max})\right]>c_{\alpha,0}\right)$
$\geq 1-P\left(\hat{F}_{1,L}(\hat{t}_{max})-F_{1,L}(\hat{t}_{max})>c_{\alpha,1}\right)-P\left(F_{0,L}(\hat{t}_{max})-\hat{F}_{0,L}(\hat{t}_{max})>c_{\alpha,0}\right)$
$=1-P\left(\hat{F}_{1,L}(\hat{t}_{max})-F_{1,L}(\hat{t}_{max})>c_{\alpha,1}\right)$
$\hskip 20.0pt-P\left(\hat{F}_{0,L}(\hat{t}_{max})-F_{0,L}(\hat{t}_{max})>c_{\alpha,0}\right)$ (A.1)
$\geq 1-P\left(\sup_{t}\left(\hat{F}_{1,L}(t)-F_{1,L}(t)\right)>c_{\alpha,1}\right)-P\left(\sup_{t}\left(\hat{F}_{0,L}(t)-F_{0,L}(t)\right)>c_{\alpha,0}\right)$
$=1-\mathbb{E}\left[P\left(\sup_{t}\left(\hat{F}_{1,L}(t)-F_{1,L}(t)\right)>c_{\alpha,1}\Bigm{|}\mathcal{I}_{A}\right)\right]$
$\hskip 15.0pt-\mathbb{E}\left[P\left(\sup_{t}\left(\hat{F}_{0,L}(t)-F_{0,L}(t)\right)>c_{\alpha,0}\Bigm{|}\mathcal{I}_{A}\right)\right]$ (A.2)
$\geq 1-\mathbb{E}\left[\exp{(-2|\mathcal{I}^{M}_{1}|c_{\alpha,1}^{2})}\right]-\mathbb{E}\left[\exp{(-2|\mathcal{I}^{M}_{0}|c_{\alpha,0}^{2})}\right]$ (A.3)
$=1-\exp{(-2|\mathcal{I}^{M}_{1}|c_{\alpha,1}^{2})}-\exp{(-2|\mathcal{I}^{M}_{0}|c_{\alpha,0}^{2})}$
$=1-\alpha$
(A.1) holds since $\hat{F}_{0,L}(t)-F_{0,L}(t)$ is symmetric around zero for all $t$. (A.2) holds by the law of iterated expectations, and note the expected values are in respect to $\mathcal{I}_{A}$. (A.3) holds by the one-sided DKW inequality as in Massart (1990). Note the use of sample splitting and conditioning on the auxiliary set $\mathcal{I}_{A}$ is important since it ensures that $(Y_{i},\hat{s}_{L}(X_{i}),D_{i})_{i\in\mathcal{I}^{1}_{M}\cup\mathcal{I}^{0}_{M}}$ are iid, assumption required for the DKW inequality.
The proof for $P\left(\theta\in\left[0,\widetilde{\theta}_{U}+c_{\alpha}\right]\right)\geq 1-\alpha$ is analogous. Finally, the validity of the two-sided confidence interval follows immediately:
$P\Biggl{(}\theta\in\left[\widetilde{\theta}_{L}-c_{\alpha/2},\widetilde{\theta}_{U}+c_{\alpha/2}\right]\Biggr{)}$ $=P\Biggl{(}\theta\in\left[\widetilde{\theta}_{L}-c_{\alpha/2},1\right]\land\theta\in\left[0,\widetilde{\theta}_{U}+c_{\alpha/2}\right]\Biggr{)}$
$=1-P\Biggl{(}\theta\not\in\left[\widetilde{\theta}_{L}-c_{\alpha/2},1\right]\lor\theta\not\in\left[0,\widetilde{\theta}_{U}+c_{\alpha/2}\right]\Biggr{)}$
$\geq 1-P\Biggl{(}\theta\not\in\left[\widetilde{\theta}_{L}-c_{\alpha/2},1\right]\Biggr{)}-P\Biggl{(}\theta\not\in\left[0,\widetilde{\theta}_{U}+c_{\alpha/2}\right]\Biggr{)}$
$\geq 1-\alpha/2-\alpha/2$
$=1-\alpha$
Definitions used in the proofs of the results of Section 4:
  • $t_{L,P}=t_{max,P},\ t_{U,P}=t_{min,P}$;
  • $F_{P,j}(t,s)=P(Y(j)\leq s(X)+t)$ (note $F_{P,j}(t,s)$ is random if $t$ and/or $s$ are random);
  • $\mathcal{I}_{j,k}=\{i\in\{1,\dots,n\}:D_{i}=j\land i\in\mathcal{I}_{k}\}$;
  • $\widehat{F}_{j,k}(t,s)=\frac{1}{n_{j}/K}\sum_{i\in\mathcal{I}_{j,k}}\mathbb{I}(Y_{i}\leq s(X_{i})+t)$;
  • $\hat{t}_{L}=\arg\max_{t}\frac{1}{K}\sum_{k=1}^{K}\left[\widehat{F}_{1,k}(t,\hat{s}_{L,k})-\widehat{F}_{0,k}(t,\hat{s}_{L,k})\right]$;
  • $\hat{t}_{U}=\arg\min_{t}\frac{1}{K}\sum_{k=1}^{K}\left[\widehat{F}_{1,k}(t,\hat{s}_{U,k})-\widehat{F}_{0,k}(t,\hat{s}_{U,k})\right]$;
  • $\hat{\pi}=n_{1}/n$.
Theorem 3 is restated as Theorem A.1:
Theorem A.1 (Asymptotic distribution of cross-fitting estimators).
Let $\mathcal{P}$ denote a set of probability functions that satisfy Assumptions 1 and 2. Let $\{P_{n}\}_{n\geq 1}\subset\mathcal{P}$ be a sequence of probability functions such that
$\begin{pmatrix}\sigma^{2}_{L,P_{n}}&\sigma_{L,U,P_{n}}\\ \sigma_{L,U,P_{n}}&\sigma^{2}_{U,P_{n}}\end{pmatrix}\to\begin{pmatrix}\sigma^{2}_{L}&\sigma_{L,U}\\ \sigma_{L,U}&\sigma^{2}_{U}\end{pmatrix}$
for some $\sigma^{2}_{L},\sigma^{2}_{U}\geq 0$ and $\sigma_{L,U}\in\mathbb{R}$.
Then, there exists $\bar{\theta}_{L,P_{n}},\bar{\theta}_{U,P_{n}}$ with $\bar{\theta}_{L,P_{n}}\leq\theta_{L,P_{n}}^{*}\leq\theta_{P_{n}}\leq\theta_{U,P_{n}}^{*}\leq\bar{\theta}_{U,P_{n}}$ such that
[…]
Proof of Theorem A.1.
For $A\in\{L,U\}$, define
  • $\bar{\theta}_{A,P,k}=F_{P,1}(\hat{t}_{A},\hat{s}_{A,k})-F_{P,0}(\hat{t}_{A},\hat{s}_{A,k})$,
  • $\bar{\theta}_{A,P}=\frac{1}{K}\sum_{k=1}^{K}\bar{\theta}_{A,P,k}$,
  • $\widehat{\theta}_{A,k}=\widehat{F}_{1,k}(\hat{t}_{A},\hat{s}_{A,k})-\widehat{F}_{0,k}(\hat{t}_{A},\hat{s}_{A,k})$.
Note that $\widehat{\theta}_{A}=\frac{1}{K}\sum_{k=1}^{K}\widehat{\theta}_{A,k}$, and that
$\widehat{\theta}_{A,k}-\bar{\theta}_{A,P_{n},k}=\left[\widehat{F}_{1,k}(\hat{t}_{A},\hat{s}_{A,k})-F_{P_{n},1}(\hat{t}_{A},\hat{s}_{A,k})\right]-\left[\widehat{F}_{0,k}(\hat{t}_{A},\hat{s}_{A,k})-F_{P_{n},0}(\hat{t}_{A},\hat{s}_{A,k})\right]$
The proof consists in showing, for $A\in\{L,U\}$ and $j=0,1$, that
$\sqrt{\frac{n_{j}}{K}}\left(\widehat{F}_{j,k}(\hat{t}_{A},\hat{s}_{A,k})-F_{P_{n},j}(\hat{t}_{A},\hat{s}_{A,k})\right)$
$=\sqrt{\frac{n_{j}}{K}}\left(\widehat{F}_{j,k}(\hat{t}_{A},s_{A,P_{n}})-F_{P_{n},j}(\hat{t}_{A},s_{A,P_{n}})\right)+o_{P_{n}}(1)$ (A.4)
$=\sqrt{\frac{n_{j}}{K}}\left(\widehat{F}_{j,k}(t_{A,P_{n}},s_{A,P_{n}})-F_{P_{n},j}(t_{A,P_{n}},s_{A,P_{n}})\right)+o_{P_{n}}(1)$ (A.5)
Finally, I show
[…]
Note $\bar{\theta}_{L,P_{n}}\leq\theta_{L,P_{n}}^{*}\leq\theta_{P_{n}}\leq\theta_{U,P_{n}}^{*}\leq\bar{\theta}_{U,P_{n}}$ holds since, for each $k$, $\theta_{P_{n}}\in[\bar{\theta}_{L,P_{n},k},\bar{\theta}_{U,P_{n},k}]$ holds from (2.3).
  • Step 1: Equation (A.4)
    I show that
    $\left(\sqrt{\frac{n_{j}}{K}}\right)^{-1}\sum_{i\in\mathcal{I}_{j,k}}\Bigl{[}\left(\mathbb{I}(Y_{i}\leq\hat{s}_{A,k}(X_{i})+\hat{t}_{A})-\mathbb{I}(Y_{i}\leq s_{A,P_{n}}(X_{i})+\hat{t}_{A})\right)$
    $\hskip 90.0pt-\left(F_{P_{n},j}(\hat{t}_{A},\hat{s}_{A,k})-F_{P_{n},j}(\hat{t}_{A},s_{A,P_{n}})\right)\Bigr{]}$
    $=:\left(\sqrt{\frac{n_{j}}{K}}\right)^{-1}\sum_{i\in\mathcal{I}_{j,k}}W_{i}(\hat{t}_{A})$
    $\leq\sup_{t}\left(\sqrt{\frac{n_{j}}{K}}\right)^{-1}\sum_{i\in\mathcal{I}_{j,k}}W_{i}(t)=o_{P_{n}}(1)$
    The result follows from stochastic equicontinuity of $\left(\sqrt{\frac{n_{j}}{K}}\right)^{-1}\sum_{i\in\mathcal{I}_{j,k}}W_{i}(t)$ plus pointwise convergence in $t$ to zero (see, e.g., Theorem 22.9 in Davidson (2021)).
    Note that
    $\left(\sqrt{\frac{n_{j}}{K}}\right)^{-1}\sum_{i\in\mathcal{I}_{j,k}}\Bigl{[}\left(\mathbb{I}(Y_{i}\leq\hat{s}_{A,k}(X_{i})+t)-\mathbb{I}(Y_{i}\leq s_{A,P_{n}}(X_{i})+t)\right)$
    $\hskip 90.0pt-\left(F_{P_{n},j}(t,\hat{s}_{A,k})-F_{P_{n},j}(t,s_{A,P_{n}})\right)\Bigr{]}$
    $=\left(\sqrt{\frac{n_{j}}{K}}\right)^{-1}\sum_{i\in\mathcal{I}_{j,k}}\left(\mathbb{I}(Y_{i}\leq\hat{s}_{A,k}(X_{i})+t)-F_{P_{n},j}(t,\hat{s}_{A,k})\right)$
    $\hskip 20.0pt-\left(\sqrt{\frac{n_{j}}{K}}\right)^{-1}\sum_{i\in\mathcal{I}_{j,k}}\left(\mathbb{I}(Y_{i}\leq s_{A,P_{n}}(X_{i})+t)-F_{P_{n},j}(t,s_{A,P_{n}})\right),$
    and so it is enough to show stochastic equicontinuity of
    $\left(\sqrt{\frac{n_{j}}{K}}\right)^{-1}\sum_{i\in\mathcal{I}_{j,k}}\left(\mathbb{I}(Y_{i}\leq\tilde{s}(X_{i})+t)-F_{P_{n},j}(t,\tilde{s})\right)$
    for an arbitrary random variable $\tilde{s}$ independent of $\mathcal{I}_{j,k}$. This follows since =$\mathcal{F}=\{f_{t}:\mathbb{R}\rightarrow[0,1],w\to f_{t}(w)=\mathbb{I}(w\leq t):t\in\mathbb{R}\}$ is Donsker and Pre-Gaussian uniformly in $P\in\{P^{*}:(Y-s(X),D,X)\sim P^{*},(Y,D,X)\sim P^{\prime}\in\mathcal{P},s\in\{s^{\prime}:\mathcal{X}\to\mathcal{Y}\}\}$. This is the case since the covering number of $\mathcal{F}$ for any such $P$ does not depend on $s$.
    Finally, for verifying pointwise convergence, note that
    $\mathbb{E}[W_{i}(t)]=\mathbb{E}[W_{i}(t)|\hat{s}_{A,k}]=0.$
    Hence, it is enough to show that
    $Var\left[\left(\sqrt{\frac{n_{j}}{K}}\right)^{-1}\sum_{i\in\mathcal{I}_{j,k}}W_{i}(t)\right]$ $=\mathbb{E}\left[Var\left[\left(\sqrt{\frac{n_{j}}{K}}\right)^{-1}\sum_{i\in\mathcal{I}_{j,k}}W_{i}(t)\Bigm{|}\hat{s}_{A,k}\right]\right]$
    $=\mathbb{E}\left[Var\left[W_{i}(t)\Bigm{|}\hat{s}_{A,k}\right]\right]$
    $\leq Var\left[W(t)\right]\to 0$
    where the first equality follows from the Law of Total Variance, and the second from the fact that $\{W_{i}(t)\}_{i\in\mathcal{I}_{j,k}}$ are iid conditional on $\hat{s}_{A,k}$ for any $t$. We have
    $Var\left[W(t)\right]$
    $\leq\mathbb{E}\left[W(t)^{2}\right]$
    $\leq 2\mathbb{E}\left[\left|\mathbb{I}(Y_{i}\leq\hat{s}_{A,k}(X_{i})+t)-\mathbb{I}(Y_{i}\leq s_{A,P_{n}}(X_{i})+t)\right|\right]$
    $\hskip 5.0pt+2\mathbb{E}\left[\left(F_{P_{n},j}(t,\hat{s}_{A,k})-F_{P_{n},j}(t,s_{A,P_{n}})\right)^{2}\right]$
    $\leq 4\left|F_{P_{n},j}(t,\hat{s}_{A,k})-F_{P_{n},j}(t,s_{A,P_{n}})\right|,$
    and $\left|F_{P_{n},j}(t,\hat{s}_{A,k})-F_{P_{n},j}(t,s_{A,P_{n}})\right|\to 0$ follows from equicontinuity of $P(Y(j)\leq t)$ by A2(i) and the fact that $|\hat{s}_{A,k}(X)-s_{A,P_{n}}(X)|\overset{P_{n}}{\longrightarrow}0$ by A2(ii).
  • Step 2: Equation (A.5)
    The result follows from (see, for example, Lemma 19.24 in van der Vaart (1998))
    • (i)
      Asymptotic equicontinuity of
      $\mathbb{G}_{A,P,j,k}(t)=\sqrt{\frac{n_{j}}{K}}\left(\widehat{F}_{j,k}(t,s_{A,P_{n}})-F_{P_{n},j}(t,s_{A,P_{n}})\right)$
      uniformly in $P\in\mathcal{P}$;
    • (ii)
      $|\hat{t}_{A}-t_{A,P_{n}}|\overset{P_{n}}{\longrightarrow}0$;
    • (iii)
      $\mathbb{E}\left[\left|\mathbb{I}(Y(j)\leq s_{A,P_{n}}+\hat{t}_{A})-\mathbb{I}(Y(j)\leq s_{A,P_{n}}+t_{A,P_{n}})\right|^{2}\right]\to 0$, where the expectation takes $\hat{t}_{A}$ fixed.
    (i) follows from the same argument for stochastic equicontinuity as in Step 1 (equation A.4). (ii) follows from consistency of M-estimators (e.g., Corollary 3.2.3 in Van Der Vaart and Wellner (2023)), given that $P(Y(j)\leq t)$ is equicontinuous by A2(i) and $t_{A,P}$ is unique and well-separated by A2(iii); (iii) follows from equicontinuity of $P(Y(j)\leq t)$ by A2(i) and since $|\hat{t}_{A}-t_{A,P_{n}}|\overset{P_{n}}{\longrightarrow}0$.
  • Step 3: Equation (A.6)
    We have that
    $\sqrt{n}(\widehat{\theta}_{A}-\bar{\theta}_{A,P_{n}})$
    $=\sqrt{n}\frac{1}{K}\sum_{k=1}^{K}\left(\widehat{\theta}_{A,k}-\bar{\theta}_{A,P_{n},k}\right)$
    $=\sqrt{n}\frac{1}{K}\sum_{k=1}^{K}\Bigl{(}\left[\widehat{F}_{1,k}(\hat{t}_{A},\hat{s}_{A,k})-F_{P_{n},1}(\hat{t}_{A},\hat{s}_{A,k})\right]$
    $\hskip 70.0pt-\left[\widehat{F}_{0,k}(\hat{t}_{A},\hat{s}_{A,k})-F_{P_{n},0}(\hat{t}_{A},\hat{s}_{A,k})\right]\Bigr{)}$
    $=\sqrt{n}\frac{1}{K}\sum_{k=1}^{K}\Bigl{(}\left[\widehat{F}_{1,k}(t_{A,P_{n}},s_{A,P_{n}})-F_{P_{n},1}(t_{A,P_{n}},s_{A,P_{n}})\right]$
    $\hskip 70.0pt-\left[\widehat{F}_{0,k}(t_{A,P_{n}},s_{A,P_{n}})-F_{P_{n},0}(t_{A,P_{n}},s_{A,P_{n}})\right]\Bigr{)}$
    $\hskip 15.0pt+\sqrt{n}\frac{1}{K}\sum_{k=1}^{K}\left(\frac{\sqrt{K}}{\sqrt{n_{1}}}o_{P_{n}}(1)-\frac{\sqrt{K}}{\sqrt{n_{0}}}o_{P_{n}}(1)\right)$ (A.7)
    $=\frac{1}{\sqrt{\hat{\pi}}}\frac{1}{\sqrt{n_{1}}}\sum_{i:D_{i}=1}\left[\mathbb{I}(Y_{i}\leq s_{A,P_{n}}(X_{i})+t_{A,P_{n}})-F_{P_{n},1}(t_{A,P_{n}},s_{A,P_{n}})\right]$
    $\hskip 15.0pt-\frac{1}{\sqrt{1-\hat{\pi}}}\frac{1}{\sqrt{n_{0}}}\sum_{i:D_{i}=0}\left[\mathbb{I}(Y_{i}\leq s_{A,P_{n}}(X_{i})+t_{A,P_{n}})-F_{P_{n},0}(t_{A,P_{n}},s_{A,P_{n}})\right]$
    $\hskip 15.0pt+o_{P_{n}}(1)$
    where (A.7) follows from (A.4) and (A.5).
    Finally, the result follows from a triangular array Central Limit Theorem (e.g., Theorems 27.2 and 27.3 in Billingsley (1995)), with joint normality following from the Cramér-Wold device (e.g., p. 18 in Serfling (1980)).
Proof of Theorem 4.
The inequalities for the one-sided confidence intervals follow directly from consistency of $(\hat{\sigma}_{L},\hat{\sigma}_{U})$ and from Theorem A.1, since $\bar{\theta}_{L,P_{n}}\leq\theta_{L,P_{n}}^{*}\leq\theta_{P_{n}}\leq\theta_{U,P_{n}}^{*}\leq\bar{\theta}_{U,P_{n}}$. For the two-sided case, the inequality follows from Theorem A.1 and consistency of $(\hat{\sigma}_{L},\hat{\sigma}_{U},\hat{\sigma}_{L,U})$ by applying Proposition 3 in Stoye (2009) centering the estimators at the outer bounds $\bar{\theta}_{L,P_{n}}$ and $\bar{\theta}_{U,P_{n}}$. Since these are in general not sharp, the equality in Stoye (2009) becomes an inequality.
Consistency of $(\hat{\sigma}_{L},\hat{\sigma}_{U},\hat{\sigma}_{L,U})$ follows from the weak law of large numbers, equicontinuity of $P(Y(j)\leq t)$ by A2(i), $|\hat{s}_{A,k}(X)-s_{A,P_{n}}(X)|\overset{P_{n}}{\longrightarrow}0$ by A2(ii) and $\hat{t}_{A}-t_{A,P_{n}}\overset{P_{n}}{\longrightarrow}0$ (see Step 2 in proof of Theorem A.1). ∎
Proof of Theorem 5.
[…]
$\bar{\theta}_{A,P,k}-\theta_{A,P}^{*}=[F_{P,1}(\hat{t}_{A},\hat{s}_{A,k})-F_{P,0}(\hat{t}_{A},\hat{s}_{A,k})]-[F_{P,1}(t_{A,P},s^{*}_{A,P})-F_{P,0}(t_{A,P},s^{*}_{A,P})],$
[…]
To show $\sqrt{n}(\bar{\theta}_{A,P_{n}}-\theta_{A,P_{n}}^{*})=o_{P_{n}}(1)$, it is enough to show $\sqrt{n}\left(\bar{\theta}_{A,P_{n},k}-\theta_{A,P_{n}}^{*}\right)=o_{P_{n}}(1)$ since $K$ is fixed. Define $\Delta_{P}(t,s|x)=F_{P,1}(t,s|x)-F_{P,0}(t,s|x)$, where $F_{P,j}(t,s|x)=P(Y(j)\leq s(X)+t|X=x)$ and note that $\Delta_{P}(t,s)=\mathbb{E}[\Delta_{P}(t,s|X)]$. Then, $|\bar{\theta}_{A,P_{n},k}-\theta_{A,P_{n}}^{*}|=|\Delta_{P_{n}}(\hat{t}_{A},\hat{s}_{A,k})-\Delta_{P_{n}}(0,s^{*}_{A,P_{n}})|$, since $t_{A,P_{n}}=0$ by definition of $s^{*}_{A,P_{n}}$ (see proof of Theorem 1).
It also holds that $|\Delta_{P_{n}}(\hat{t}_{A},\hat{s}_{A,k})-\Delta_{P_{n}}(0,s^{*}_{A,P_{n}})|\leq|\Delta_{P_{n}}(0,\hat{s}_{A,k})-\Delta_{P_{n}}(0,s^{*}_{A,P_{n}})|$. To see this, focus on the case of the lower bound (the upper bound case is analogous). We have from Theorem 1 that $\Delta_{P_{n}}(0,s^{*}_{L,P_{n}})\geq\Delta_{P_{n}}(\hat{t}_{L},\hat{s}_{L,k})$, and by definition of $\hat{t}_{L}$ it holds that $\Delta_{P_{n}}(\hat{t}_{L},\hat{s}_{L,k})\geq\Delta_{P_{n}}(0,\hat{s}_{L,k})$. Hence, $|\Delta_{P_{n}}(\hat{t}_{L},\hat{s}_{L,k})-\Delta_{P_{n}}(0,s^{*}_{L,P_{n}})|=\Delta_{P_{n}}(0,s^{*}_{L,P_{n}})-\Delta_{P_{n}}(\hat{t}_{L},\hat{s}_{L,k})\leq\Delta_{P_{n}}(0,s^{*}_{L,P_{n}})-\Delta_{P_{n}}(0,\hat{s}_{L,k})$. Therefore, it is enough to show that $\sqrt{n}|\Delta_{P_{n}}(0,\hat{s}_{A,k})-\Delta_{P_{n}}(0,s^{*}_{A,P_{n}})|=o_{P_{n}}(1)$.
A Taylor expansion gives
$\Delta_{P_{n}}(0,s|x)$ $=\Delta_{P_{n}}(0,s^{*}_{A,P_{n}}|x)+\Delta^{\prime\prime}_{P_{n}}(0,\tilde{s}_{A,P_{n}}|x)(s(x)-s^{*}_{A,P_{n}}(x))^{2}$
where $\tilde{s}_{A,P_{n}}(x)$ is in between $s(x)$ and $s^{*}_{A,P_{n}}(x)$, and $\Delta^{\prime}_{P_{n}}(0,s^{*}_{A,P_{n}}|x)=0$ by definition of $s^{*}_{A,P_{n}}$. Hence,
$|\Delta_{P_{n}}(0,s|x)-\Delta_{P_{n}}(0,s^{*}_{A,P_{n}}|x)|\lesssim(s(x)-s^{*}_{A,P_{n}}(x))^{2}$ (A.8)
where $\lesssim$ means bounded above up to a universal constant (not depending on $n$, $P_{n}$, $t$ or $s$), since the second derivative is uniformly bounded by assumption. Hence, we have
$\sqrt{n}\left(\bar{\theta}_{A,P_{n},k}-\theta_{A,P_{n}}^{*}\right)$ $=\sqrt{n}\left[\Delta_{P_{n}}(\hat{t}_{A},\hat{s}_{A,k})-\Delta_{P_{n}}(0,s^{*}_{A,P_{n}})\right]$
$\leq\sqrt{n}\mathbb{E}\left[\left|\Delta_{P_{n}}(0,\hat{s}_{A,k}|X)-\Delta_{P_{n}}(0,s^{*}_{A,P_{n}}|X)\right|\right]$
$\lesssim\sqrt{n}\mathbb{E}\left[\left(\hat{s}_{A,k}(X)-s^{*}_{A,P_{n}}(X)\right)^{2}\right]$
which is $o_{P_{n}}(1)$ by assumption. ∎
Online Appendix OA
Supplemental Appendix to “Predicting the Distribution of Treatment Effects via Covariate-Adjustment,
with an Application to Microcredit”
OA.1 Relaxing constant propensity score
I propose two alternatives to the cross-fitting estimators in Definition 2(iii). For simplicity, I focus on the lower bound.
  • (A)
    Constant propensity score within groups: Let there be fixed $\bar{G}>0$ groups, such that the propensity score is fixed within-group (and bounded away from zero and one). Define $n_{1,g}$ as the number of observations in group $g$ assigned to treatment and $n_{0,g}$ the number of observations assigned to control. Finally, let $G_{i}\in\{1,\dots,\bar{G}\}$ be the group to which observation $i$ belongs. Then, a new cross-fitting estimator for the lower bound can defined as
    $\widehat{\theta}^{A}_{L}$ $=\max_{t}\Biggl{\{}\frac{1}{\bar{G}}\sum_{g=1}^{\bar{G}}\Biggl{[}\frac{1}{n_{1,g}}\sum_{\begin{subarray}{c}i:D_{i}=1\\ G_{i}=g\end{subarray}}\mathbb{I}(Y_{i}\leq\hat{s}_{L,k(i)}(X_{i})+t)$
    $\hskip 60.0pt-\frac{1}{n_{0,g}}\sum_{\begin{subarray}{c}i:D_{i}=0\\ G_{i}=g\end{subarray}}\mathbb{I}(Y_{i}\leq\hat{s}_{L,k(i)}(X_{i})+t)\Biggr{]}\Biggr{\}},$
    where the data is split such that for each fold $k$, there is balance in the number of treated and control units within each group. The asymptotic distribution follows very similarly to the proof of Theorem 3 by applying the arguments in (A.4) and (A.5) by each group-fold combination, instead of by fold. The asymptotic variance-covariance matrix of the estimators is given by
    $\sigma^{2}_{A,n}=\sum_{g=1}^{\bar{G}}\pi(g)^{2}\left(\frac{1}{\pi_{1}(g)}Var[Z_{A}(1)|G=g]+\frac{1}{\pi_{0}(g)}Var[Z_{A}(0)|G=g]\right),$
    $\sigma_{L,U,n}=$
    $\sum_{g=1}^{G}\pi(g)^{2}\left(\frac{1}{\pi_{1}(g)}Cov[Z_{L}(1),Z_{U}(1)|G=g]+\frac{1}{\pi_{0}(g)}Cov[Z_{L}(0),Z_{U}(0)|G=g]\right),$
    where $A\in\{L,U\}$, and
    $\pi(g)=P(G=g),\quad\pi_{j}(g)=P(D=j|G=g),$
    $Z_{A}(j)=\mathbb{I}(Y(j)\leq s_{A,P}(X)+t_{A,P}).$
    As before, the covariance matrix can be estimated with sample analogues.
  • (B)
    Known propensity score: Assume the propensity score $p(x)=P(D=1|X=x)$ is known and bounded away from zero and one. Then, the estimator for the lower bound can be defined as
    $\widehat{\theta}^{B}_{L}$ $=\max_{t}\Biggl{\{}\frac{1}{n}\sum_{i=1}^{n}\biggl{(}\frac{D_{i}}{p(X_{i})}\mathbb{I}(Y_{i}\leq\hat{s}_{L,k(i)}(X_{i})+t)$
    $\hskip 60.0pt-\frac{1-D_{i}}{1-p(X_{i})}\mathbb{I}(Y_{i}\leq\hat{s}_{L,k(i)}(X_{i})+t)\biggr{)}\Biggr{\}}.$
    The proof of Theorem 3 extends to this case with minor adjustments. Note that since $p(x)$ is (uniformly) bounded away from zero and one, the same Donsker properties hold for $(\widehat{\theta}^{B}_{L},\widehat{\theta}^{B}_{U})$. The asymptotic variance-covariance matrix of the estimators is given by
    $\sigma^{2}_{A,P}=\mathbb{E}_{P}\left[\left(\frac{D}{p(X)^{2}}+\frac{1-D}{(1-p(X))^{2}}\right)\mathbb{I}(Y\leq s_{A,P}(X)+t_{A,P})\right]-\theta_{A,P}^{2},$
    $\sigma_{L,U,P}=$
    $\mathbb{E}_{P}\left[\left(\frac{D}{p(X)^{2}}+\frac{1-D}{(1-p(X))^{2}}\right)\mathbb{I}(Y\leq s_{L,P}(X)+t_{L,P})\mathbb{I}(Y\leq s_{U,P}(X)+t_{U,P})\right]$
    $-\theta_{L,P}\theta_{U,P},$
    where $A\in\{L,U\}$ and $p(x)=P(D=1|X=x)$.
OA.2 Definition of Two-Sided Confidence Intervals
The definition of $\widehat{CI}_{\alpha}$ follows from Stoye (2009). Let $h_{n}$ denote a preassigned sequence such that $h_{n}\to 0$ and $\sqrt{n}h_{n}\to\infty$, and define
$\Lambda=\begin{cases}\widehat{\theta}_{U}-\widehat{\theta}_{L},&\widehat{\theta}_{U}-\widehat{\theta}_{L}>h_{n}\\ 0,&\text{otherwise}\end{cases}$
Then, the CI is defined by
$\widehat{CI}_{\alpha}=\begin{cases}\left[\widehat{\theta}_{L}-\widehat{c}_{L}\frac{\widehat{\sigma}_{L}}{\sqrt{n}},\widehat{\theta}_{U}+\widehat{c}_{U}\frac{\widehat{\sigma}_{U}}{\sqrt{n}}\right],&\text{ if }\widehat{\theta}_{L}-\widehat{c}_{L}\frac{\widehat{\sigma}_{L}}{\sqrt{n}}\leq\widehat{\theta}_{U}+\widehat{c}_{U}\frac{\widehat{\sigma}_{U}}{\sqrt{n}}\\ \emptyset,&\text{ otherwise}\end{cases}$
where $(\widehat{c}_{L},\widehat{c}_{U})$ minimize $(\widehat{\sigma}_{L}\widehat{c}_{L}+\widehat{\sigma}_{U}\widehat{c}_{U})$ subject to the constraints that
$P\left(-\widehat{c}_{L}\leq Z_{1}\land\frac{\widehat{\sigma}_{L,U}}{\widehat{\sigma}_{L}\widehat{\sigma}_{U}}Z_{1}\leq\widehat{c}_{U}+\frac{\sqrt{n}\Lambda}{\widehat{\sigma}_{U}}+Z_{2}\sqrt{1-\frac{\widehat{\sigma}^{2}_{L,U}}{\widehat{\sigma}^{2}_{L}\widehat{\sigma}^{2}_{U}}}\right)$ $\geq 1-\alpha$
$P\left(-\widehat{c}_{L}-\frac{\sqrt{n}\Lambda}{\widehat{\sigma}_{L}}+Z_{2}\sqrt{1-\frac{\widehat{\sigma}^{2}_{L,U}}{\widehat{\sigma}^{2}_{L}\widehat{\sigma}^{2}_{U}}}\leq\frac{\widehat{\sigma}_{L,U}}{\widehat{\sigma}_{L}\widehat{\sigma}_{U}}Z_{1}\land Z_{1}\leq\widehat{c}_{U}\right)$ $\geq 1-\alpha$
and $Z_{1}$, $Z_{2}$ are independent standard normal random variables.
Importantly, calculating $\widehat{CI}_{\alpha}$ relies on choosing the hyperparameter $h_{n}$. A possible choice is Stoye’s suggestion of $h_{n}=n^{-1/2}(\log\log n)^{1/2}$, and other possibilities are $h_{n}=n^{-1/2}(\log n)^{1/2}$ and $h_{n}=n^{-1/2}(Q\log\log n)^{1/2}$ for $Q\geq 2$, as suggested in Andrews and Soares (2010). Discussing the best choice for $h_{n}$ is beyond the scope of this paper, but a practical procedure for applications is to assess the sensitivity of the CI to different choices of $h_{n}$.
OA.3 Relaxing Assumption 2(iii)
I propose two alternatives to Assumption 2(iii). The first requires changing the definition of the estimators $\widehat{\theta}_{L}$ and $\widehat{\theta}_{U}$, while the second relaxes the assumption by studying the asymptotic distribution of the original estimators without the condition of a unique maximizer.
  • (A)
    Changing the estimators: The first alternative is to change the definition of $\widehat{\theta}_{L}$ and $\widehat{\theta}_{U}$ to avoid the optimization problem. For example, $t$ could be fixed at $t=0$ or learned from other folds inside the cross-fitting procedure. Focusing on the case of $\widehat{\theta}_{L}$, when estimating $\hat{s}_{L,k}$, one could estimate $\hat{t}_{k}$ (with data from all except the $k$-th fold) which maximizes the expression inside the $k$-th fold, and estimate the lower bound as
    $\frac{1}{n_{1}}\sum_{i:D_{i}=1}\mathbb{I}(Y_{i}\leq\hat{s}_{L,k(i)}(X_{i})+\hat{t}_{k})-\frac{1}{n_{0}}\sum_{i:D_{i}=0}\mathbb{I}(Y_{i}\leq\hat{s}_{L,k(i)}(X_{i})+\hat{t}_{k})$
    This would be valid since it is equivalent to a new estimator $\tilde{s}_{L,k}=\hat{s}_{L,k}+\hat{t}_{k}$, and there is no optimization in the final estimator.
  • (B)
    Asymptotic distribution without unique maximizer: For simplicity, I focus on the case of the lower bound. Define
    $\widehat{\Delta}(t)=\frac{1}{n_{1}}\sum_{i:D_{i}=1}\mathbb{I}(Y_{i}\leq\hat{s}_{L,k(i)}(X_{i})+t)-\frac{1}{n_{0}}\sum_{i:D_{i}=0}\mathbb{I}(Y_{i}\leq\hat{s}_{L,k(i)}(X_{i})+t),$
    and note that $\widehat{\theta}_{L}=\max_{t}\widehat{\Delta}(t)$. From the proof of Theorem A.1 (eq. A.5), it follows that
    $\mathbb{G}_{n,k}(t)$ $=\sqrt{n}\left(\widehat{\Delta}(t,\hat{s}_{L,k})-\Delta_{P}(t,\hat{s}_{L,k})\right)$
    $=\sqrt{n}\left(\widehat{\Delta}(t,s_{L,P})-\Delta_{P}(t,s_{L,P})\right)+o_{P}(1)$
    converges to a Gaussian process. Note that $\sup_{t}\Delta_{P}(t,s_{L,P})\leq\theta^{*}_{L,P}$ (equation 2.3). If $\Delta_{P}(t,s_{L,P})\neq 0$ for at least one value of $t$, the $\sup$ function is Hadamard differentiable at $\Delta_{P}(\cdot,s_{L,P})$ (Cárcamo et al., 2020). It follows that
[…]
    uniformly in $P\in\mathcal{P}$, where $M=\{t:\Delta_{P}(t,s_{L,P})=\sup_{t^{\prime}}\Delta_{P}(t^{\prime},s_{L,P})\}$ and $\mathbb{G}_{P}$ is a Gaussian process with covariance function given by
    $\frac{Cov[Z_{t}(1),Z_{t^{\prime}}(1)]}{\pi}+\frac{Cov[Z_{t}(0),Z_{t^{\prime}}(0)]}{1-\pi},$
    where $Z_{t}(j)=\mathbb{I}(Y(j)\leq s_{L,P}+t)$, $\pi=P(D=1)$.
    Inference can be conducted by approximating the distribution of $\sup_{t\in M}\mathbb{G}_{P}(t)$. This could be done, for example, by conservatively estimating $M$ with $\widehat{M}$, such that $M\subseteq\widehat{M}$ with probability approaching one. Alternatively, one could apply the bootstrap. The validity of the bootstrap follows by combining the argument of Cárcamo et al. (2020), who show that the $\sup$ function is Hadamard differentiable, with a conditional functional delta method argument (e.g., Theorem 3.10.11 in Van Der Vaart and Wellner (2023)).
OA.4 Additional Tables
Table 4: Bounds on fraction harmed by microcredit - alternative values of $\delta$
Lower Bound
($\delta=-0.05$)
Upper Bound
($\delta=0.05$)
Bosnia
0.053
[0.027]
0.971
[0.054]
Egypt
0.182
[0.000]
0.749
[0.000]
Philippines (1)
0.034
[0.181]
0.950
[0.101]
Philippines (2)
0.082
[0.006]
0.853
[0.000]
South Africa
0.022
[0.130]
0.873
[0.000]
  • Standard errors in parentheses, and p-values in brackets. The null hypothesis for the lower bound is $\theta^{*}_{L}=0$, and for the upper bound $\theta^{*}_{U}=1$. Response variable: $log(1+household\ income)$.
[ht] Lower Bound Lower Bound (No Covariates) Upper Bound Upper Bound (No Covariates) Bosnia 0.012 [0.354] 0.006 [0.171] 0.918 [0.003] 0.936 [0.007] Egypt 0.249 [0.000] 0.008 [0.328] 0.828 [0.000] 0.958 [0.008] Philippines (1) 0.077 [0.025] 0.071 [0.036] 0.902 [0.006] 0.931 [0.005] Philippines (2) 0.094 [0.002] 0.046 [0.085] 0.961 [0.005] 0.981 [0.089] All 0.108 [0.000] 0.033 [0.000] 0.902 [0.000] 0.951 [0.000]
Table 5: Bounds on fraction harmed by microcredit - business profits
  • p-values in brackets. The null hypothesis for the lower bound is $\theta^{*}_{L}=0$, and for the upper bound $\theta^{*}_{U}=1$. $1-$ Upper Bound represents fraction benefited by microcredit.